### v1

# JMA: a General Algorithm to Craft Nearly Optimal Targeted Adversarial Examples

Abstract
Most of the approaches proposed so far to craft targeted adversarial examples against Deep Learning classifiers are highly suboptimal and typically rely on increasing the likelihood of the target class, thus implicitly focusing on one-hot encoding settings. In this paper, we propose a more general, theoretically sound, targeted attack that resorts to the minimization of a Jacobian-induced MAhalanobis distance (JMA) term, taking into account the effort (in the input space) required to move the latent space representation of the input sample in a given direction. The minimization is solved by exploiting the Wolfe duality theorem, reducing the problem to the solution of a Non-Negative Least Square (NNLS) problem. The proposed algorithm provides an optimal solution to a linearized version of the adversarial example problem originally introduced by Szegedy et al. [1]. The experiments we carried out confirm the generality of the proposed attack which is proven to be effective under a wide variety of output encoding schemes. Noticeably, the JMA attack is also effective in a multi-label classification scenario, being capable to induce a targeted modification of up to half the labels in a complex multilabel classification scenario with 20 labels, a capability that is out of reach of all the attacks proposed so far. As a further advantage, the JMA attack usually requires very few iterations, thus resulting more efficient than existing methods.
Index Terms:
Adversarial Examples, Deep Learning Security, Adversarial Machine Learning, Multi-Label Classification, Mahalanobis Distance, Non-Negative Least Square Problems
publicationid: pubid: $0$–$00$ 2021 IEEE
1 Introduction
Adversarial examples, namely, quasi-imperceptible perturbations capable to induce an incorrect decision, are a serious threat to deep-learning classifiers [1, 2]. Since the publication of the seminal work by Szegedy et al. [1] in which the existence of adversarial examples was first observed, a large number of gradient-based methods have been proposed to implement adversarial attacks against Deep Neural Networks (DNNs) in white-box and black-box scenarios. Most of the attacks proposed so far rely on greedy algorithms that permit to find an effective adversarial example in a reasonable amount of time [2, 3, 4, 5, 6, 7].
Most attacks create the adversarial examples by minimizing a cost function subject to a constraint on the maximum perturbation introduced in the image. Moreover, they focus on classifiers based on one-hot encoding. In this setting, the final activation layer consists in the application of an activation function (usually a softmax) and a normalization, mapping the last layer outputs, called logits, into a probability vector, associating a probability value to each class. The loss function corresponds to the categorical cross entropy. As a consequence, in the targeted case, adversarial examples focus on raising the probability value of the target class. This is obviously the best strategy with one-hot encoding in the presence of a softmax function, however, this strategy is no more optimal in the presence of output encoding schemes based on channel coding, and in the case of multi-label classification.
A more flexible attack working at the logits level has been proposed by Carlini and Wagner [7]. The attack works by decreasing the difference between the largest logit and the logit of the target class (in the targeted case). Working at the logits level allows to avoid vanishing gradient problems [8], hence Carlini and Wagner method often yields better performance compared to methods that directly minimize the loss term. However, working on two logits at a time, without considering the effect of the perturbation on the other logits, is clearly suboptimum in general, with the consequence that the approach in [7] often results in a greedy, lengthy, iterative minimization process. The problem is more evident when the attacker aims at attacking a DNN adopting different encoding mechanisms, like Error Correction Output Coding (ECOC) [9], and also multi-label DNNs. In this case, all the logits have to be modified simultaneously.
Crafting adversarial examples for networks adopting a generic output encoding scheme, and for multi-label classification, is a very challenging task, that, to the best of our knowledge, has not been given much attention, with the exception of some scattered works proposing suboptimum approaches [10, 11]. From the attacker side, the difficulties of applying an optimal attack in this case are due to the fact that the labels are not mutually exclusive (like in the one- hot case), and decreasing or increasing the network output in correspondence of some nodes may change the values taken by the other nodes in an unpredictable way.
To overcome the above drawbacks, we introduce a new, theoretically sound, targeted adversarial attack, named Jacobian-induced Mahalanobis distance Attack (JMA). JMA solves the original formulation of the adversarial example problem introduced by Szegedy et al. [1], which aims at minimizing the strength of the adversarial perturbation subject to the constraint that the image is classified as belonging to the target class. The solution is found via a two-step procedure: i) given a target point in the output space, the optimum perturbation moving the input sample to the target point is first determined, under a linear assumption on the effect of the distortion on the model output; ii) then, the target point that minimizes the perturbation introduced by the attack is determined by minimizing the Mahalanobis distance induced by the Jacobian matrix of the network input-output function. By exploiting the Wolfe duality theorem, we reduce the problem to the solution of a Non-Negative Least Square (NNLS) problem, that can be efficiently solved numerically. The optimal perturbation derived in this way is then applied to the input image. Ideally, the JMA attack should produce the desired adversarial image in just one-shot. However, due to the local optimality of the proposed approach, in practice, it is sometimes necessary to carry out some iterations to obtain a valid adversarial example, recomputing the Jacobian matrix every time.
The experiments we carried out confirm that the proposed attack is more efficient than state-of-the-art attacks from a computational perspective, requiring a much lower number of iterations to carry out the attack. State-of-the-art attacks are also outperformed in terms of distortion and attack success rate. The effectiveness of the proposed adversarial attack method is particularly evident when it is used to attack networks adopting Error Correction Output Coding (ECOC) [9] and in multi-label classification scenarios, where the JMA attack is capable to simultaneously change in a desired way up to 10 out of 20 labels of the output label vector, a capability which is out of reach of the algorithms proposed so far. We also verified that the method has a very good behavior in the one-hot encoding scenario, where it achieves performance that are comparable to the state-of-the-art with a significant reduction of the computing time. In summary, the main contributions of the paper are:
  • We introduce a new targeted adversarial attack, named Jacobian-induced Mahalanobis distance Attack (JMA). The attack resorts to the minimization of a Jacobian-induced MAhalanobis distance (JMA) term, with the Jacobian matrix taking into account the effort (in the input space) required to move the latent space representation of the input sample in a given direction. In theory, the algorithm is one-shot.
  • We solve the constrained minimization of the Mahalanobis distance by exploiting the Wolfe duality theorem, reducing the problem to the solution of a non-negative least square (NNLS) problem, that can be solved numerically.
  • We validate the new attack considering several datasets (CIFAR-10, GTSRB, MNIST, VOC2012) and networks adopting different encoding schemes, namely the ECOC scheme, multi-label classification, and also one-hot encoding. Our experiments confirm that the proposed attack is very general and can work in all these cases, being very efficient, often requiring only very few iterations to attack the image.
  • We compare JMA against several state-of-the-art methods, showing that the proposed attack is more efficient and requires a much lower number of iterations. In the cases of ECOC and multi-label classification, it also achieves lower distortion and higher attack success rate. In particular, in the multi-label classification scenario, the proposed attack is capable to change up to half labels of the sample output vector, even when the number of labels increases.
The rest of the paper is organized as follows: in Section 2, we provide the main formalism and describe the various classification frameworks addressed in the paper. In Section 3, we review the most relevant adversarial attacks introduced so far, with particular attention to the attacks against DNNs adopting with output channel encoding. Then, in Section 4, we present the JMA adversarial attack. The experimental methodology and setting are described in Section 5. Sections 6 and 7 report and discuss the results of the experiments. The papers ends in Section 8 with our final conclusions.
2 Background and Notation
In this section, we review the main output encoding schemes for DNN classifiers.
Let $x\in\mathbb{R}^{m}$ denote the input of the network, and $y(x)$ (or simply $y$) the class $x$ belongs to. We indicate with $l$ the number of classes. To classify $x$, the neural network first maps $x$ into a reduced dimensional space $\mathbb{R}^{n}$ ($n<m$). Every class is associated to an output label column vector $c_{k}=(c_{k1},c_{k2},...,c_{kn})^{T}$. Let $f(x):\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ be a column vector indicating the end-to-end neural network function, and $f_{i}(x)$ the $i$-th element of the vector. Classification in favour of one of the $l$ classes is obtained by applying a function $\phi$ to $f(x)$, the exact form of $\phi$ depending on the output encoding scheme used by the network. In the following, we denote by $z=(z_{1},z_{2},...,z_{n})^{T}$ the vector with the logit values, that is, the values of the network nodes before the final activation function, which is responsible to map the output of the penultimate layer of the network into the [0,1] (sometimes [-1, 1]) range. Given an image $x$, the notation $f^{-j}(x)$ is used to refer to the internal model representation at layer $L-j$ ($L$ being the total number of layers of the network). With this notation, $z(x)=f^{-1}(x)$.
2.1 One-hot encoding
In the case of networks adopting the one-hot encoding scheme, the number of output nodes corresponds to the number of classes ($l=n$), and $c_{k}$ is a binary vector with all 0’s except for position $k$, where it takes value 1. In this case, the length of $z$ is $l$, and the logits are directly mapped onto the output nodes through a softmax function as follows:
$f_{k}(x)=\frac{\exp(z_{k})}{\sum_{i=1}^{l}\exp(z_{i})},$ (1)
for $k=1,..,l$. This allows to interpret $f_{k}(x)$ as the probability assigned to class $k$, and the final prediction is made by letting $\phi(x)={\arg\max}_{k}\hskip 2.84544ptf_{k}(x)$.
Typically, in the one-hot encoding case, training is carried out by minimizing the categorical cross-entropy loss defined as $\mathcal{L}(x,y)=-\sum_{i=1}^{l}c_{yi}\log(f_{i}(x))=-\log f_{y}(x)$ [^1] We remind that $y$ denotes the class sample $x$ belongs to..
2.2 Error-correction-output-coding (ECOC)
Sometimes the output classes are encoded by using the codewords of a channel code. In this way, error correction can be applied to recover from incorrect network behaviors. In this case, the network output dimension $n$ corresponds to the length of the codewords and $c_{k}$ is the codeword associated to the $k$-th output class. The number of classes is typically less than $2^{n}$ ($l<2^{n}$). An example of the use of channel coding to define the class label vectors is given in [12] (ECOC). The use of ECOC has also been proposed as a way to improve the robustness against adversarial attacks in a white-box setting [9]. The rationale is the following: while with classifiers based on standard one-hot encoding the attacker can induce an error by modifying one single logit (the one associated to the target class), the final decision of the ECOC classifier depends on multiple logits in a complicated manner, and hence it is supposedly more difficult to attack.
Formally, with ECOC, a distinct codeword $c_{k}$ is assigned to every output class. Let $C=\{c_{1},c_{2},\cdots,c_{l}\}$ define the codebook, that is, the matrix of codewords. Each element of $C$ can take values in $\{-1,1\}$. To compute the output of the network the logits are mapped into the $[-1,1]$ range by means of an activation function $\sigma()$, that is $f(x)=\sigma(z)$, where $\sigma()$ is applied element-wise to the logits. A common choice for $\sigma()$ is the $\tanh$ function. To make the final decision, the probability of class $k$ is first computed:
$p_{k}(x)=\frac{\max(f(x)^{T}c_{k},0)}{\sum_{i=1}^{l}\max(f(x)^{T}c_{i},0)},$ (2)
where $f(x)^{T}c_{k}$ is the inner product between $f(x)$ and $c_{k}$. Since $c_{ij}$’s take values in $\{-1,1\}$, the max is necessary to avoid negative probabilities. Then, the model’s final prediction is given by $\phi(x)={\arg\max}_{k}\hskip 2.84544ptp_{k}(x)$. Note that Eq. (2) reduces to Eq. (1) in the case of one-hot encoding, when ${C}=\mathbf{I}_{l}$ and where $\mathbf{I}_{l}$ is the identity $l\times l$ matrix.
In the ECOC case, training is usually carried out by minimizing the hinge loss function, defined as $\mathcal{L}(x,y)=\sum_{i=1}^{n}\max(1-c_{yi}f_{i}(x),0)$.
2.3 Multi-label classification
In this paper, we refer to multi-label classification as a scenario wherein the classifier is asked to decide about the presence of absence within the image of $n$ image characteristics or features. For example, the classifier may be asked to detect the presence of $n$ possible classes of objects, or decide about $n$ binary attributes (like indoor/outdoor, night/day, sunny/rainy). The presence/absence of the looked-for features is encoded by the $n$ outputs of the network (see, for instance, [13, 14]). In general, the number of possible outputs is $l=2^{n}$, and each output is encoded in a matrix $C$ having size $n\times 2^{n}$. In fact, it is possible that some labels’ combinations are not feasible. In such a case, $C$ contains only the allowed combinations, somewhat playing the same role of channel coding. In this paper, by multi-label classification, we indicate a situation where all combinations of labels are possible.
In the case of multi-label classification, given the logit vector $z$, the activation function $\sigma()$ is applied element-wise to the components of $z$, and the prediction on each label component is made component-wise. Assuming that $c_{k}\in\{0,1\}^{n}$ for any $k$, and that the logistic function is applied to the logits, decoding corresponds to 0.5-thresholding each output score independently [^2] If $c_{k}\in\{-1,1\}^{n}$, the tanh activation is applied followed by 0-thresholding..
For the loss function, a common choice is the binary cross-entropy loss $\mathcal{L}(x,y)=-\sum_{i=1}^{n}(c_{yi}\log(f_{i}(x))+(1-c_{yi})\log(1-f_{i}(x)))$.
3 Adversarial attacks against DNNs
The vulnerability of deep neural networks (DNNs) to adversarial samples has been first pointed out by Szegedy et al. in [1]. When the goal of the attacker is to induce a generic misclassification, we talk about an untargeted attack. When the misclassification is in favor of a specific class, we say that the attack is a targeted one. According to [1], the generation of a targeted adversarial example can be formally described as
$&\textrm{minimize}\quad||\delta||_{2}\\ &\textrm{s.t.}\quad\phi(x+\delta)=t\neq\phi(x)\\ &\textrm{and}\quad x+\delta\in[0,1]^{m}\end{aligned},$ (3)
where $||\delta||_{2}$ denotes the $l_{2}$ norm of the perturbation $\delta$, and $t$ is the target class of the attack. The same formulation holds for the untargeted case, with the first constraint replaced by $\phi(x+\delta)\neq\phi(x)$. The constraint $x+\delta\in[0,1]^{m}$ makes sure that the resulting adversarial example is a valid input.
3.1 Basic adversarial attacks
Solving the minimization in (3) is generally hard, then the adversarial examples are determined by solving the simplified problem described below:
$\min_{\delta}$ $\hskip 5.69046pt\mathcal{L}({x+\delta},t)+\lambda\cdot\left\|\delta\right\|_{2}^{2}$ (4)
s.t. $\hskip 5.69046ptx+\delta\in[0,1]^{m}$
where $\mathcal{L}({x+\delta},t)$ is the loss function under the target class $t$, and $\lambda$ is a parameter balancing the importance of the two terms. The problem in (4) can be solved by the box-constraint L-BFGS method [1]. Specifically, a line search is carried out to find the value of $\lambda>0$ for which the solution of (4) satisfies the adversarial condition, that is $\phi(x+\delta)=t$. Usually, the categorical cross entropy loss is considered. From (4), it is immediate to see that, in the one-hot case, when the categorical cross entropy loss is considered, only the target class contributes to the loss term, and then the attack only cares about increasing $f_{t}(x)$, regardless of the other scores, all the more that, due to the presence of the softmax, this also implies decreasing the other outputs.
To reduce the complexity of the L-BFGS attack, several suboptimal solutions have been proposed, that consider the problem - equivalent to (4) - of minimizing the loss function $\mathcal{L}$ subject to a constraint on the perturbation $\delta$. The Fast Sign Gradient Method (FGSM) method [2] obtains an adversarial perturbation in a computationally efficient way by considering the sign of the gradient of the output with respect to the input image. The FGSM attack is a one-step attack whose strength can be modulated only by increasing the value of the term multiplying the gradient sign. In [4], an iterative version of FGSM is introduced by applying FGSM multiple times, with a smaller step size, each time by recomputing the gradient. This method is often referred to as I-FGSM, or Basic Iterative Method (BIM). Another attack similar to the FGSM attack is the projected gradient descent (PGD) attack [5], that can be regarded to as a multi-step extension of the FGSM attack where the clip operation performed by BIM on the gradient (to force the solution to stay in the [0,1] range) is replaced by the gradient projection.
Other gradient-based attacks are the Jacobian-based Saliency Map Attack (JSMA) [3] and the DeepFool attack [6]. JSMA [3] consists of a greedy iterative procedure which relies on forward propagation to compute, at each iteration, a saliency map, indicating the pixels that contribute most to the classification. In such attack, the Jacobian matrix of the network function is used to compute the forward derivative and build the saliency map. A variant of this attack, called Maximal JSMA, has been proposed in [15]. DeepFool [6] is an efficient iterative attack that considers the minimal perturbation with respect to a linearized classifier. In this way, the current version of the adversarial example is projected onto the separating affine hyperplane at each iteration, until the classification boundary is crossed.
The above gradient-based methods often fail due to the gradient vanishing problem [7]. A solution to overcome this issue has been proposed in [16], where the values of the logits are scaled by a temperature parameter. However, this approach lacks flexibility and the suitable distillation temperature parameter strongly depends on the network. A more flexible and efficient method working at the logits level has been proposed by Carlini and Wagner (C&W) [7]. In the C&W attack the loss term is replaced by a properly designed function depending on the difference between the logit of the target class and the largest among the logits of the other classes (since the logits do not sum one, two logits at a time must be considered). The perturbation is kept in the valid range by properly modifying the objective function. In this way, the following unconstrained problem is considered and solved (for the case of $L_{2}$ distortion metric):
$\min_{w}$ $\hskip 5.69046pt||\delta(w)||_{2}^{2}+$
$\lambda\cdot\max(\max_{i\neq t}z_{i}(x+\delta(w))-z_{t}(x+\delta(w)),-\kappa),$ (5)
where $\delta(w)=1/2(\tanh(w)+1)-x$ is the distortion introduced in the image, and the constant parameter $\kappa>0$ is used to encourage the attack to generate a high confidence attacked image. Since logits are more sensitive to modifications of the input than the probability values obtained after the softmax activation C&W attack is less sensitive to vanishing gradient problems. However, we notice that, by considering only two logits at time, and neglecting the effect of the perturbation on the other logits, this approach is particularly suited to the one-hot case and is highly suboptimum when other output encoding schemes are used.
3.2 Attacks against DNNs with output encoding
The adversarial attack algorithms described so far focus on models based on one-hot encoding. This is a favorable scenario for the attacker, who only needs to focus on increasing the output score (or logit) of the target class $t$. This is not the case with DNNs based on different output encoding schemes. The simplest examples are multi-label classifiers. From the attacker’s side, the difficulties of carrying out the attack in this case are due to the fact that the labels are not mutually exclusive, and decreasing or increasing the network output in correspondence of some nodes may change the values taken by the other nodes in an unpredictable way. In addition, the attacker may want to modify more than one output label in a desired way. For instance, he may want to induce the classifier to interpret a daylight image showing a car driving in the rain, as a night image of a car driving with no rain. This prevents the application of the basic adversarial algorithms described in the previous section.
The problem of crafting adversarial examples against multi-label classifiers has been addressed in [10], where the C&W and the DeepFool attacks have been extended to work in a multi-label setting. The extension of C&W attack, referred to as ML-C&W, works as follows: for every output node, a hinge loss term similar to the one in Eq. (3.1) is considered in the minimization. More specifically, the following term is added to the minimization: $\sum_{i=0}^{n}\max(0,\gamma-c_{ti}\cdot z_{i}(x+\delta))$, where $c_{t}$ is the target label vector and $\gamma$ is a confidence parameter. Note that the extension implicitly assumes that decoding is carried out element-wise. This is not true when all labels’ combinations are possible and with schemes adopting output channel coding. However, by working at the logits level, the C&W attack is general and can also be extended to work with different output encoding schemes (as discussed below).
An extension of the DeepFool attack, referred to as ML-DF, has also been proposed in [10]. Instead of targeting an objective function minimizing the distortion, the method looks for the minimum perturbation that allows to enter the target region under a linear assumption on the model behavior. ML-DF works at the output score level, and hence is more prone to the vanishing gradient problem than ML-C&W. Given the threshold vector ${\rho}=({\rho}_{1},{\rho}_{2},..,{\rho}_{l})^{T}$ for the score level output ($\rho_{i}=0.5$ when $f_{i}\in[0,1]$, 00 when $f_{i}\in[-1,1]$), a vector with the distance to the boundary is derived and used to compute the perturbation. The perturbation obtained in this way is applied to the image. ML-DF applies the above procedure iteratively until the attack succeeds or a maximum number of iterations is reached. ML-DF is a greedy method and often leads to very large distortions; Moreover, the performance obtained are always inferior with respect to ML-C&W [10, 11].
Inspired by ML-DF, another approach to implement an adversarial attack in the multi-label case, named Multi-Label Attack via Linear programming (MLA-LP), has been proposed in [11]. By assuming that the loss changes are linear for small distortions, a simplified formulation of the attack is solved to minimize the $L_{\infty}$ distortion introduced in the image subject to a constraint on the loss value, requiring that the final loss is lower than the threshold value. Thanks to the adoption of the $L_{\infty}$ distortion, the problem can be easily transformed into a linear programming problem, for which several solvers exist. Similarly to ML-DF, MLA-LP works at the final score output level. By considering the loss function instead of the model output function, MLA-LP tends to modify less the confidences of non-attacked labels with respect to ML-DF, with a lower distortion for the resulting attacked image.
When the output coding scheme is based on ECOC, an adversarial attack can be carried out by extending to this setting the C&W method. Specifically, the C&W attack can be applied at the probability level, exploiting the specific mapping of the logits to probabilities. In particular, C&W attack can be used to attack networks adopting ECOC output encoding by replacing the logit terms $z_{i}(x)$ with $f(x)^{T}c_{i}$, that is, by considering the loss term $\max(\max_{i\neq t}f(x+\delta(w))^{T}c_{i}-f(x+\delta(w))^{T}c_{t},-\kappa)$.
In the case of an ECOC-based network, adversarial examples can also be obtained by using the attack described in [17] (in the following, this method is referred to as ECOC attack). Such an attack incorporates within the minimization problem the ECOC decoding procedure. Formally, the optimization problem solved in [17] is defined as:
$\min_{\delta}$ $\hskip 5.69046pt\left(||\delta||_{2}^{2}-\lambda\cdot\min_{i}(c_{ti}\cdot z_{i}(x+\delta),\eta)\right),$ (6)
where $\eta$ is a constant parameter setting a confidence threshold for the attack.
Although not specialized for this case, ML-C&W can also be used to attack ECOC-based networks. Instead, ML-DeepFool and MLA-LP, that work at the score output level and requires knowledge of the threshold vector $\rho$, cannot be extended to the case of networks adopting an output encoding scheme like ECOC. In fact, the threshold vector $\rho$ is not available in the ECOC case since the decision is made after the correlations are computed. A straightforward, highly suboptimum, way to apply these methods to the ECOC case is by setting $\rho=(0,0,...,0)^{T}$, as if all the codewords were possible, and performing the decoding element-wise.
3.2.1 The LOTS attack
A completely different kind of attack, which can be applied to any network regardless of the output encoding scheme is the Layerwise Origin-Target Synthesis (LOTS) attack [18]. LOTS generates the adversarial example by modifying the input sample so that its representation in the feature space is as similar as possible to that of a given target sample. Formally, given the internal representation $f^{-j}(x)$ of an image $x$ at layer $L-j$ and a target internal representation $f^{-j}(x_{t})$, LOTS attempts to minimize the term:
$||f^{-j}(x^{t})-f^{-j}(x)||^{2}.$ (7)
At each iteration, the algorithm updates the input sample as
$x^{(i)}=x^{(i-1)}-\frac{\nu^{-j}(x,x^{t})}{\max_{x}(|\nu^{-j}(x,x^{t})|)}$ (8)
where $\nu^{-j}(x,x^{t})=\nabla_{x}\left(||f^{-j}(x^{t})-f^{-j}(x)||^{2}\right)$, and ||$\cdot$|| ⋅ | is applied element-wise, until the Euclidean distance between $f^{-l}(x)$ and the target is smaller than a predefined threshold. The final perturbed image $x^{adv}$ obtained in this way has an internal representation at layer $L-j$ that mimics that of the target sample $x^{t}$. In the LOTS algorithm, the target point $x^{t}$ is any point belonging to the target class of the attack. We observe that in general such $x^{t}$ may not be available, e.g. in the multilabel case when the attacker targets an arbitrary labels’ combination.
When applied to the logits level ($f^{-1}$), LOTS simultaneously modifies all the logits. However, by focusing on the minimization of the Euclidean loss, the distortion introduced in the input space is not controlled.
3.2.2 JMA and prior art
Below we discuss the relationship and advantages of JMA with respect to the state-of-the-art attacks described above.
In contrast to LOTS [18], JMA defines the target point in the feature space that minimizes perturbation necessary to move the feature representation of the input sample to the target point. It does so, by relying on the Mahalanobis distance induced by the Jacobian matrix of the neural network function so to take into account the different effort, in terms of input perturbation, required to move the sample in different directions.
The ML-DF and MLA-LP methods [10, 11] are also based on the Jacobian matrix. However, in [10], the Jacobian matrix is used to implement a greedy algorithm, while in [11] is used to solve a suboptimum formulation of the attack. In our case, instead, we are able to solve the original formulation of the attack by Szegedy et al. under a first-order approximation. With such an approximation, in fact, the solution of the attack problem can be reduced to the solution of a constrained quadratic programming problem (where the objective function is a Mahalanobis distance term induced by the Jacobian). Moreover, our algorithm is a general one and can work either at the output level or at the logits level, regardless of the output encoding scheme used by the network.
4 The JMA attack
As mentioned in the introduction, to avoid vanishing gradient problems, we look for a method that can be applied at the logits level. As we said, while modifying one or two logits at a time allows to carry out a close-to-optimal adversarial attack in the one-hot encoding case [7], in the more general case of DNNs based on output encoding, this approach is highly suboptimum. In this scenario, in addition to considering all the logits simultaneously, as done in the ML-C&W attack, the correlation among the logits and the effort required in the input space to move the input sample in a given direction (in the logit or feature space) must be considered. The JMA attack tackles the above problems and finds the optimal adversarial image, by solving the optimization problem in Eq. (3), under a first-order approximation of the behavior of the network function.
In the following, we look at the output of the neural network function $f$  [^3] For ease of notation, in the following, we consider directly the model output $f(x)$, however the same the arguments cam be applied at the logit level. as a generic point in $\mathbb{R}^{n}$. According to Eq. (3), the optimum target point for the attacker corresponds to the point in the target decision region that can be reached by introducing the minimal perturbation in the image $x$. Such a point is not necessarily the point in the target decision region closest to the current output point $f(x)$ in the Euclidean norm, since evaluating the distance in the output space does not take into account the effort necessary to move the input sample into the desired output point. Let $x_{0}$ be the to-be-attacked image and let us denote with $\delta$ the perturbation applied by the attacker. Let $r=f(x_{0})+d$. The goal of the attacker is to determine the displacement $d$, in such a way that $r$ lies inside the decision region of the target class, and for which the Euclidean norm of the perturbation $\delta(d)$ necessary to reach $r$ is minimum, that is:
$\min_{d:\phi(f(x_{0})+d)=t}\left[\min_{\begin{subarray}{c}\delta:f(x_{0}+\delta)-f(x_{0})=d\\ x_{0}+\delta\in[0,1]^{m}\end{subarray}}\hskip 5.69046pt||\delta(d)||_{2}\right].$ (11)
Without loss of generality, we will assume that all the label codewords have the same norm, and that $\phi()$ applies a minimum distance decoding rule. In this case, forcing $\phi(f(x_{0})+d)=t$ is equivalent to impose the following conditions:
$r^{T}c_{t}\geq r^{T}c_{i},\quad i=1,2\dots l,$ (12)
that can be rewritten as a function of $d$ as
$d^{T}(c_{i}-c_{t})\leq g_{ti},\quad i=1,2\dots l,$ (13)
where $g_{ti}=(f(x_{0})^{T}c_{t}-f(x_{0})^{T}c_{i})$, and hence the optimization problem in Eq. (11) can be rewritten as:
$\min_{\begin{subarray}{c}d:d^{T}(c_{i}-c_{t})\leq g_{ti}\\ \forall i,i=1,\cdots,l\end{subarray}}\bigg{[}\min_{\delta:f(x_{0}+\delta)-f(x_{0})=d}\hskip 5.69046pt||\delta(d)||_{2}\bigg{]}.$ (16)
Note that we neglected the constraint $x_{0}+\delta\in[0,1]^{m}$, trusting that if $\delta$ is small the constraint is always satisfied. We will then reconsider the effect of this constraint at the end of our derivation in Section 4.1.
Under the assumption that the input perturbation is small (which should always be the case with adversarial examples), we can consider a first order approximation of the effect of the perturbation $\delta$ on the network output:
$f(x_{0}+\delta)\simeq f(x_{0})+J_{x_{0}}\delta,$ (17)
[…]
$J_{x}=\nabla_{x}f(x)|_{x_{0}}=\left[\frac{\partial f_{i}(x_{0})}{\partial x_{j}}\right]_{i=1,\cdots,n,j=1,\cdots m}.$ (18)
We now show that, under the first order approximation in Eq. (17), we can solve the minimization in Eq. (16), by applying the following two steps: i) we first determine the optimum perturbation moving the input sample to a given target point $r$; ii) then, we find the optimum target point $r^{*}$ by solving a constrained quadratic programming problem. Addressing i) corresponds to solving the following problem:
$\delta^{*}(d)=\arg\min_{\delta:f(x+\delta)-f(x_{0})=d}||\delta||_{2},$ (19)
that, under the linear approximation, becomes [^5] To keep the notation light, here and in the sequel, we omit the subscript of the Jacobian matrix.
$\delta^{*}(d)=\arg\min_{\delta:d=J\delta}\|\delta\|_{2},$ (20)
that corresponds to finding the minimum norm solution of the system of linear equations $d=J\delta$. The solution of the above problem is given by the following proposition.
Proposition 1.
Given that $n<m$, under the assumption that $J$ has full rank, the solution of the minimization in Eq. (20) is:
${\delta}^{*}(d)=J^{T}(JJ^{T})^{-1}d.$ (21)
Proof.
Eq. (20) can be rewritten as the solution of the following non-negative least-squares problem:
$\min_{\delta}\|J\delta-d||_{2}^{2}=\min_{\delta}\left\{\delta^{T}(J^{T}J)\delta-2\delta^{T}J^{T}d+d^{T}d\right\}.$ (22)
In fact, since $(J^{T}J)$ is positive definite, the to-be-minimized function in the right-hand side of the equation is (strictly) convex and thus admits a minimum. If $J$ has full rank, the solution of the above problem is unique and is given by [19]:
$\delta^{*}=J^{T}(JJ^{T})^{-1}d,$ (23)
which satisfies $J\delta^{*}=d$ (and hence $\|J\delta^{*}-d||_{2}^{2}=0$), thus also corresponding to the minimum norm solution in (20). ∎
From Proposition 1, we can express the norm of the optimum perturbation as follows
$||\delta^{*}(d)||^{2}=$ $\delta^{*T}\delta^{*}=(J^{T}{(JJ^{T})}^{-1}d)^{T}(J^{T}{(JJ^{T})}^{-1}d)$
$=$ $d^{T}(J^{T}{(JJ^{T})}^{-1})^{T}(J^{T}{(JJ^{T})}^{-1}d)$
$=$ $d^{T}{{(JJ^{T})}^{-1}}d,$ (24)
where in the last equality we have exploited the symmetry of $(JJ^{T})^{-1}$ (being the inverse of a symmetric matrix $(JJ^{T})$), and then ${((JJ^{T})^{-1})}^{T}=(JJ^{T})^{-1}$.
Given the above, the optimal target point $r^{*}$, and hence the optimal distance term $d^{*}$, is determined by solving the following minimization:
$\min_{d:d^{T}(c_{i}-c_{t})\leq g_{ti},\forall i\neq t}$ $\hskip 5.69046ptd^{T}{{(JJ^{T})}^{-1}}d.$ (25)
Note that the above formulation corresponds to minimizing the Mahalanobis distance induced by the Jacobian matrix, between the current output of the network and the target point. The above formulation has a very intuitive meaning, which is illustrated in Fig. 1. The term $d^{T}{{(JJ^{T})}^{-1}}d$ allows to take into account the effort required in the pixel domain to move along a given direction in the output domain.
Figure 1: Illustration of the intuition behind the formulation in (25). While $r^{\prime}$ is the point in the target region closest to $f(x_{0})$ in the Euclidean norm, $r^{*}$ is the point that can be reached with the minimum distortion in the input space.
To solve (25), we find convenient to rewrite it as follows:
$\min_{d:Ad\leq b}\frac{1}{2}d^{T}{{(JJ^{T})}^{-1}}d,$ (26)
where $A=[(c_{1}-c_{t})^{T}\dots(c_{l}-c_{t})^{T}]$ has size $(l-1)\times n$, and $b$ is an $l$-long vector defined as $b_{i}=g_{ti}$. We observe that when $l>n+1$ the system is overdetermined.
Theorem 1.
Problem (26) (and hence (25)) has a unique solution given by:
$d^{*}=-(JJ^{T})(A^{T}\lambda^{*}),$ (27)
where
$\lambda^{*}=\arg\min_{\lambda}\hskip 5.69046pt\frac{1}{2}{\lambda}^{T}A(JJ^{T})A^{T}\lambda+b^{T}\lambda,\quad\lambda\geq 0,$ (28)
and then
$\delta^{*}=J^{T}(JJ^{T})^{-1}d^{*}=-J^{T}(A^{T}\lambda^{*}).$ (29)
Proof.
The optimal solution of the quadratic problem in (26) can be obtained by solving the easier Wolfe dual problem [20] defined as:
$\max_{d,\lambda}\hskip 5.69046pt$ $\frac{1}{2}d^{T}{{(JJ^{T})}^{-1}}d+\lambda^{T}(Ad-b)$
s.t. ${{(JJ^{T})}^{-1}}d+A^{T}\lambda=0$
$\lambda\geq 0,$ (30)
where $\lambda$ is a column vector with $k-1$ entries. By rewriting the objective function as $-\frac{1}{2}d^{T}{{(JJ^{T})}^{-1}}d+d^{T}({{(JJ^{T})}^{-1}}d+A^{T}\lambda)-b^{T}\lambda$ and after some algebra, the optimization problem can be rephrased as:
$\min_{d,\lambda}\hskip 5.69046pt$ $\frac{1}{2}d^{T}{{(JJ^{T})}^{-1}}d+b^{T}\lambda$
s.t. ${{(JJ^{T})}^{-1}}d+A^{T}\lambda=0$
$\lambda\geq 0.$ (31)
Solving the equality constraint yields $d^{*}$ in (27) as a function of $\lambda$. By substituting $d^{*}$ in the objective function and exploiting the symmetry of $(JJ^{T})$ we obtain (28). ∎
The problem in (28) is a NNLS problem [20, 19], for which several numerical solvers exist [21], many of them belonging to the class of active set methods, see for instance [22, 23]. In particular, an easy-to-implement algorithm, whose complexity grows linearly with the number of label vectors, named Sequential Coordinate-Wise algorithm, has been proposed in [24].
Note that in the multi-label case, the number of label vectors grows exponentially with $n$ (since $l=2^{n}$). However, in this case, the problem can be significantly simplified, as discussed in Section 4.2.
4.1 Iterative formulation of JMA
In principle, JMA should produce an adversarial example in one-shot. In practice, however, this is not always the case, due to the fact that the linear assumption of $f$ holds only in a small neighborhood of the input $x$, hence the assumption may not be met when the distortion necessary to attack the image becomes larger. In order to mitigate this problem, we consider the following iterative version of JMA:
  • When $\delta^{*}$ does not bring $x$ into the target region, that is, $\phi(x+\delta^{*})\neq t$, we update the input in the direction given by the perturbation considering a small step $\epsilon$, thus remaining in the close vicinity of $x$.
  • When the target class is reached, a binary search is performed between the perturbed sample $x+\delta^{*}$ and the original input $x$, and the adversarial example resulting in the smallest perturbation is taken as the final output of the attack.
Finally, in practice, after that the optimal perturbation $\delta^{*}$ is superimposed to $x$, a clipping operation is performed to be sure that the solution remains in the $[0,1]^{m}$ range. Regarding the full rank of the Jacobian, this assumption has been verified experimentally and discussed in Section 5.4.
The iterative JMA attack resulting from the application of the above steps is described in Algorithm 1. In the algorithm, the maximum number of iterations (updates) is set to $n_{it,max}$.
Algorithm 1 Iterative Jacobian-induced MAhalanobis distance (JMA) attack
0:    max no. of iterations $n_{it,max}$, to be attacked image $x$, target class $t$ (target label vector $c_{t}$), updating step size $\epsilon$
0:    adversarial example $x^{adv}$
1:  for  $i\in[1,n_{it,max}]$ do
2:     calculate $f(x)$
3:     calculate the Jacobian matrix by backward propagation: $J=\nabla_{x}f(x)$
4:     calculate matrix $A$ and vector $b$
5:     calculate $\lambda^{*}$ in Eq.(28) via the sequential Coordinate-Wise algorithm in [24]
6:     calculate adversarial perturbation: $\delta^{*}=-J^{T}(A^{T}\lambda^{*})$.
7:     $x=x+\delta^{*}$
8:     $x=\text{clip}(x,0,1)$
9:     if $\phi(f(x))=t$ then
10:        while $\phi(f(x))=t$ do
11:           do binary search between $x$ and $x-\delta^{*}$
12:           $x=\text{clip}(x,0,1)$
13:        end while
14:        return $x^{adv}$
15:     else
16:        $x=(x-\delta^{*})+\epsilon\frac{\delta^{*}}{||\delta^{*}||}$
17:     end if
18:  end for
4.2 Simplified formulation for the multi-label case
In most cases, the number of classes $l$, determining the number of rows of matrix $A$ in Eq. (26), is limited. For multi-label classification, however, all labels’ combinations are possible, and $l=2^{n}$. Therefore, the size of $A$ increases exponentially slowing down the attack. Luckily, in this case, the minimization problem can be significantly simplified by properly rephrasing the constraint in (12). Specifically, since all the codeword combinations are possible, the constraint forcing the solution to lye in the desired region can be rewritten element-wise as follows:
$f_{j}\cdot c_{tj}\geq 0,\quad j=1,2\dots n.$ (32)
By rewriting (32) as a function of $d$, we get
$d_{j}\cdot c_{tj}\geq-f_{j}(x_{0})\cdot c_{tj},\quad j=1,2\dots n.$ (33)
that can be rewritten in the form $Md\leq e$, where $e$ is a $n$-length vector with elements $e_{j}=f(x_{0})_{j}\cdot c_{tj}$ and $M$ is an $n\times n$ diagonal matrix with diagonal elements $M_{jj}=-c_{tj}$. Then, the minimization problem has the same form of (26), with $M$ and $e$ replacing $A$ and $b$, and Proposition 1 remains valid, with the difference that the scalar vector $\lambda$ has now dimension $n$. Therefore, the NNLS problem can be solved by means of the sequential Coordinate-Wise algorithm in [24] with complexity $O(n)$ instead of $O(2^{n})$ as in the original formulation.
5 Experimental Setting
To evaluate the performance of JMA, we have run several experiments addressing different classification scenarios, including ECOC-based classification [9] and multi-label classification. Although less significant for our study, we have also run some experiments in the one-hot encoding scenario. In all the cases, the JMA attack is applied at the logits level.
We implemented the JMA attack by using Python 3.6.9 via the Keras 2.3.1 API. We run the experiments by using an NVIDIA GeForce RTX 2080 Ti GPU. The code, as well as the trained models and all the information for the reproducibility of the experiments are publicly available at https://github.com/guowei-cn/JMA–A-General-Close-to-Optimal-Targeted-Adversarial-Attack-with-Improved-Efficiency.git.
In the following sections, we describe the classification tasks considered in the experiments, present the evaluation methodology and comparison analysis, and detail the experimental setting.
5.1 Classification networks and settings
5.1.1 One-hot encoding
To assess the performance of JMA attack in the one-hot encoding scenario, we considered traffic sign classification on the German Traffic Sign Recognition Benchmark (GTSRB) dataset [25], with a VGG16 architecture. The error rate of the trained models on clean images is 0.005.
5.1.2 ECOC-based classifier
We considered the ECOC framework for three image classification tasks, namely MNIST classification [26], CIFAR-10 classification [27], and GTSRB classification [25]. We implemented the ECOC scheme considering an ensemble of networks, each one outputting one bit [^6] According to [9], resorting to an ensemble of networks permits to achieve better robustness against attacks.. The number of branches $h$ of the ensemble is $h=10$ for MNIST and CIFAR-10, and $h=16$ for GTSRB. We used Hadamard codewords as suggested in [9]. We considered an Hadamard code with $n=16$ for MNIST and CIFAR-10. For the GTSRB case, we set $l=32$, by selecting the classes with more examples, and used an Hadamart code with $n=32$. Following the original ECOC design, we considered the VGG16 architecture [28] as baseline, with the first 6 convolutional layers shared by all the networks of the ensemble (shared bottom), and the remaining 10 layers (the last 8 convolutional layers and the 2 fully connected layers) that have been trained separately for each ensemble branch. The weights of the Image-Net pre-trained model are used for the shared bottom.
5.1.3 Multi-label classification
For the multi-label experiments, we considered the model architecture adopted in [10] that is, a standard InceptionV3 [29]. The categorical hinge loss was used to train the model, with a standard Adam optimizer with lr=$10^{-4}$ and batch size 64.
TABLE I: Types of objects in the VOC2012 dataset.
Category Object name
Person person
Animal bird, cat, cow, dog, horse, sheep
Vehicle aeroplane, bicycle, boat, bus, car, motorbike, train
Indoor bottle, chair, dining table, potted plant, sofa, tv/monitor
The model was trained on the VOC2012 dataset [30], which is a benchmark dataset adopted in several multi-class classification works [14, 31]. This dataset has been used to train models to recognize object inside scenes. Specifically, the 11,530 images of the dataset contain objects from 20 possible object classes, see Table I (the objects are classified into four categories). Then, $n=20$ and the number of possible label vectors is $2^{n}\approx 10^{6}$. All datasets are split into training, validation, and testing subsets, with proportion 6:2:2. The performance of the network in this case are measured in terms of mean average precision (mAP), that is, the mean of the average precision (AP) for all the classes, where AP is a measure of the area under the precision-recall curve. According to our experiments, the mAP of our model is 0.93, which is aligned with [10].
5.2 Comparison with the state-of-the-art
We have compared the performance of JMA with those of the most relevant state-of-the-art attacks for each classification scenario. In all the cases, we used the codes made available by the authors in their repositories.
For the multi-label case, we considered ML-C&W [10], MLA-LP [11] and LOTS [18] (we have not considered ML-DF as its performance are always inferior to ML-C&W, see [10, 11]). As described in Section 3, LOTS can be applied to any internal layer of the network. In our experiments, for a fair comparison, we have applied LOTS at the logits layer. By following [18], for every target label vector, we randomly selected 20 images with that label vector, and averaged the logits to get the target output vector (for target label vectors for which the total number of images available is lower than 20, we used all the images). We stress that, in the way it works, LOTS can not target an arbitrary labels’ combination, since examples for that combination may not be available.
For the case of ECOC classification, the comparison is carried out against C&W, which is also the attack considered in [9], and [17] (ECOC Attack) specialized to work with the ECOC model [17]. Although suboptimal in this case, we also considered ML-C&W and MLA-LP, by applying the latter as discussed in Section 3.2. However, given the complexity of the attack in the ECOC scenario and the high sub-optimality of MLA-LP, the performance of this algorithm are extremely poor, and no image can be attacked by using this method. Finally, the performance of LOTS attack are also assessed. In this case, the target output vectors are computed by averaging on 50 images (when the number of available images is lower than 50, we averaged all of them).
For the one-hot encoding scenario, we considered the original C&W algorithm [7], which is one of the best performing white-box attack working at the logits level.
5.3 Evaluation methodology
To test the attacks in the various scenarios, we proceeded as follows: we randomly picked 200 images from the test set (among those that are correctly classified by the network model); then, for each image, we randomly picked a target label vector, different from the true label vector.
For multi-label classification, in principle, there can be - and in fact there are - combinations of labels that are not represented in the training set. it is also possible that some of these combinations do not correspond to valid label vectors, as some labels may not be simultaneously 1 (or 0). Since LOTS requires the availability of samples belonging to the target label vector, for ease of comparison, in our experiments, we considered label vectors that appear in the training set as target label vectors. For the other methods, some tests have also been carried out in the more challenging case where the target label vectors correspond to randomly chosen target vectors. Such target vectors are obtained by randomly changing a prescribed number of bits in the label vector of the to-be-attacked image.
An attack is considered successful only when the predicted label vector and the target label vector coincide. If only a subset of the labels can be modified by the attack, and the target label vector is not reached, we mark this as a failure. We measured the performance of the attacks in terms of Attack Success Rate (ASR), namely, the percentage of generated adversarial examples that are assigned to the target class. To measure the quality of the adversarial image, we considered the Peak Signal-Noise Ratio (PSNR), characterizing the distortion introduced by the attack. The PSNR is evaluated as $20\log_{10}(255\cdot\sqrt{H\times W}/||\delta||_{2})$, where $H\times W$ is the size of the image [^7] Given that the attacks are multi-step and iterative, and different iterations are necessary for the various algorithms, we can not use $\epsilon$ as an indication of the distortion.. For the tests with random choice of the target vector (for a prescribed number of flipped labels), we also report the average percentage of labels/bits successfully modified by the attack, indicated as bASR.
The computational complexity is measured by providing the average number of iterations required by the attack, and the time - in seconds - necessary to run it. In the following, we let $n_{it}$ denote the number of iterations necessary to run an attack, or the number of attack updates. The average number of iterations is denoted by $\bar{n}_{it}$.
5.4 Setting the Attack Parameters
The tunable parameters of JMA are the maximum number of iterations (${n}_{it,max}$) and the step size $\epsilon$. The iteration parameter takes also into account the number of ’for’ loops and the number of steps of the binary search in the final iteration (see Algorithm 1), the total number of updates of JMA, then, is ${n}_{it}=(v-1)+n_{bs}$, where $v$ is the number of ’for’ loops of the algorithm ($v<{n}_{it,max}$, see Algorithm 1) and $n_{bs}$ is the number of steps of the binary search. $n_{bs}$ is set to 6 in our the experiments.
For C&W, ML-C&W and ECOC attack, the parameters are the initial $\lambda$, the number of steps of the binary search $n_{bs}$, and the maximum number of iterations $n_{it,max}$. All these attacks carry out a binary search over the parameters $\lambda$ at every iteration to find the best value. Then, $n_{it,max}$ provides a constraint on $n_{bs}\cdot b$, where $b$ is the number of iterations of the optimization algorithm. As for their setting, we consider $n_{bs}=5$ and $10$, and various values of $\lambda$.
Finally, for MLA-LP and LOTS, the only parameter is the maximum number of iterations $n_{it,max}$ allowed in the gradient descent. In our experiments, we considered various values of $n_{it,max}$. For MLA-LP, we set $n_{it,max}=100$ for the multi-label classification case, and raised it up to $3000$ in the ECOC case when the attack is more difficult. For LOTS, we set $n_{it,max}=2000$, namely the maximum number considered in [18], maximizing the chances to find an adversarial example. Given that the time complexity of LOTS is limited compared to the other comparison methods, in fact, the time complexity of the attack remains small also for large values of $n_{it,max}$. We have verified experimentally that LOTS converges within 1000 iterations 92% of the times [^8] Convergence is determined by checking whether the new loss value is close enough to the average loss value of the last 10 iterations., validating the soundness of our choice. For ML-C&W, we set $n_{it,max}=1000$ as default value, and in most of the cases results are reported for the setting $(\lambda,n_{bs},n_{it,max})=(0.01,10,1000)$, that resulted in the highest ASR. Finally, for JMA, we generally got no benefit by increasing the number of iterations beyond 200, since most of the time JMA converges in few iterations, and 200 is more than enough to get and adversarial image in all the cases.
For sake of simplicity and comparison, the confidence parameter is set to 0 for all the attack methods. To avoid that the an attack gets stuck with some images (e.g. in the multi-label case when many labels have to be changed), we set a limit on the running time of the attack, and considered the attack unsuccessful when an adversarial example could not be generated in less than 1 hour.
Regarding JMA, we found experimentally that, the assumption made in the theoretical analysis that the Jacobian matrix has full rank always holds when the number of iterations of the attack remains small. However, when the number of iterations increases it occasionally happens that the rank of the Jacobian matrix is not full. In this case, an adversarial perturbation can not be found, resulting in a failure of the attack.
6 Results
The results we obtained in the various classification settings are reported and discussed in this section. In all the tables, $\bar{n}_{it}$ refers to the average number of iterations/updating of the attack, averaged on the successful attacks. The attack parameters we report are $(\lambda,n_{bs},n_{it,max})$ for C&W, ML-C&W and ECOC attack, and $(\epsilon,n_{it,max})$ for JMA.
6.1 One-hot encoding
As we said, the main advantage of JMA with respect to state of the art methods is obtained for classifiers that do not adopt one-hot encoding. In such a scenario, in fact, focusing only on the target logit or final score, and ignoring the others, as done by the most common adversarial attacks, is approximately optimal. Nevertheless, we verified that the JMA method is also effective in the one-hot encoding scenario. Table II reports the results we got on the GTSRB dataset. We see that JMA achieves a much higher ASR compared to C&W, getting close to 1, yet with a lower PSNR. This is likely due to the fact that JMA tends to modify all the output scores, while C&W directly focuses on raising the target output, which is an effective strategy when the one-hot scheme is adopted for the encoding. In addition, JMA can obtain an adversarial image in a much smaller number of iterations, requiring only very few seconds (2.45 sec in the best setting) to attack an image thus resulting much faster than C&W attack.
TABLE II: Results of the attacks against an one-hot encoding classifier for the GTSRB dataset.
Parameters ASR PSNR $\bar{n}_{it}$ Time(s)
JMA ( 0.01, 200) 0.45 43.16 66.49 9.82
(0.05, 200) 0.81 37.88 50.65 7.44
(0.1, 200) 0.88 36.86 38.00 5.66
(0.2, 200) 0.94 35.76 29.63 4.41
(0.6,200) 0.98 33.67 16.20 2.46
C&W (1e-4,5, 100) 0.19 46.65 85.95 45.06
(1e-4,5,200) 0.33 45.24 171.88 90.11
(1e-2,5,500 ) 0.55 42.36 443.18 233.45
(1e-1,10,2000) 0.75 41.14 1661.11 875.35
(1e-1,10,5000) 0.82 41.02 3577.5 1884.59
6.2 ECOC-based classifiers
The results we got for ECOC-based classification on GTSRB, CIFAR-10, and MNIST are reported in Tables III-V. We can see that, in all the cases, JMA achieves a much higher ASR for a similar PSNR, with a significantly lower computational complexity.
In the case of GTSRB (Table III), the JMA attack achieves ASR=0.98 at PSNR=38.08dB with a total average number of iterations $\bar{n}_{it}=29.15$ and attack time of $9.41$ sec. This represents a dramatic improvement with respect to the other methods. In the case of ECOC attack, for instance, the best result is ASR=0.93, reached for a similar distortion (PSNR=37.24dB), with $\bar{n}_{it}$ larger than 1000 ($\bar{n}_{it}=1089.03$), which is two orders of magnitude larger than JMA. We also see that the C&W method fails in this case, with the ASR being about 0.30 for comparable PSNR values. We remind that the C&W method is not designed to work in the ECOC case (the results are aligned with those reported in [9, 17]), and when applied to the ECOC classification scenario it looses part of its effectiveness (see discussion in Section 3). ML-C&W is more effective, however it can only achieve ASR=0.78 with a significantly lower PSNR (30.13dB) and at the price of a larger complexity. A very poor behavior of the attack is observed for LOTS and in particular MLA-LP, that is never able to attack the test images, notwithstanding the large value of $n_{it,max}$. Since this method implements a greedy approach, it is not surprising that it fails in the complex ECOC classification scenario (more in general, the performance of MLA-LP are poor whenever the attack aims at changing several bits/labels - see the results on multi-label classification).
For CIFAR-10 (Table IV), JMA achieves ASR = 1 with a PSNR of about 40dB, with very low complexity. C&W attack achieves a much lower ASR for a similar PSNR, with higher complexity. The ECOC attack can achieve ASR = 0.99 with PSNR=37.27dB in the setting with $n_{it,max}=2000$, in which case the computational complexity of the algorithm is high ($\bar{n}_{it}=1347.42$ for an average attack time of $557.86$ sec). Regarding LOTS, it can achieve ASR = 1, yet with a lower PSNR (35.79 dB). The results we got with LOTS confirm the lower complexity of this method with respect to ECOC attack and C&W. We also observe that, in this case, ML-C&W have poorer performance than C&W, with only 56% of the images that can be attacked with a lower PSNR, and with a very high computational cost. Finally, as before, MLA-LP is not able to attack the image in almost all the cases (ASR = 0.04).
The effectiveness of JMA is also verified in the MNIST case (see Table V). As it is always the case when attacking MNIST, the PSNR of the attacked images is lower since, unavoidably, the attacker has to modify more the image (this is particulary true in the targeted attack scenario). In this case, JMA reaches ASR = 1 with PSNR = 23.30 dB, while the ASR of ECOC attack (in the best possible setting) and LOTS is respectively 0.73 and 0.78 for a similar PSNR. The computational time is higher for LOTS and much higher for the ECOC attack. Both C&W and ML-C&W have poor performance and the ASR is lower than 0.52 for a PSNR similar to that obtained by JMA.
Fig. 2 reports some examples of images (successfully) attacked with the various methods for the three tasks.
Figure 2: Examples of attacked images in the ECOC classification case. The attacker’s goal is to cause a misclassification from ‘cat’ to ‘airplane’ (CIFAR-10), from speed limitation ‘80kph’ to ‘30kph’ (GTSRB), and from digit ‘1’ to ‘3’ (MNIST). The corresponding PSNR and time required for the attack are also provided for each case.
TABLE III: Results against ECOC-based classification on the GTSRB dataset.
Parameters ASR PSNR ${\scriptsize\bar{n}_{it}}$ Time(s)
JMA (0.05, 200) 0.93 39.12 58.46 18.14
(0.1, 200) 0.98 38.08 29.15 9.41
(0.2, 200) 0.99 36.98 25.30 8.27
C&W (1e-4,5,100) 0.08 49.38 65.69 51.31
(1e-4,5,200) 0.11 48.02 97.73 97.67
(1e-4,5,500) 0.12 47.42 148.33 235.76
(1e-2,5,500) 0.21 42.12 194.19 221.63
(1e-1,10,2000) 0.31 39.77 782.29 1453.85
ECOC attack (1e-4,5,100) 0.35 43.42 71.22 87.87
(1e-4,5,200) 0.40 44.18 109.90 305.28
(1e-4,5,500) 0.43 44.45 175.67 332.88
(1e-2,5,500) 0.61 40.64 309.03 501.80
(1e-1,10,2000) 0.93 37.24 1089.03 2312.64
ML-CW (0.01, 10, 1000) 0.78 30.13 111.02 2973.17
LOTS $n_{it,max}=2000$ 0.40 39.21 395.33 97.52
MLA-LP $n_{it,max}=1000$ 0 NA NA NA ($>$ 3600)
TABLE IV: Results against ECOC-based classification on CIFAR-10 dataset.
Parameters ASR PSNR $\bar{n}_{it}$ Time(s)
JMA (0.05, 200) 1.00 39.39 14.78 3.14
(0.1, 200) 1.00 38.79 6.88 1.92
(0.5, 200) 1.00 38.20 6.10 1.20
C&W (1e-4,5,100) 0.49 40.82 40.76 23.20
(1e-4,5,500) 0.60 41.42 119.45 68.49
(1e-4,10,200) 0.90 37.41 83.97 48.15
(1e-4,10,500) 0.96 37.62 187.96 107.78
(1e-1,10,2000) 0.99 37.87 593.09 340.09
ECOC attack (1e-4,5,100) 0.65 39.11 83.08 50.36
(1e-4,5,500) 0.92 38.56 337.53 180.37
(1e-4,10,200) 0.93 37.78 176.85 252.89
(1e-4,10,500) 0.95 37.33 383.74 557.86
(1e-1,10,2000) 0.99 37.27 1347.42 2173.31
ML-CW (0.01, 10, 1000) 0.565 29.86 49.57 1409.02
LOTS $n_{it,max}=2000$ 1.00 35.79 38.85 8.61
MLA-LP $n_{it,max}=1000$ 0.04 39.39 5 7.04
TABLE V: Results against ECOC-based classification on MNIST dataset.
Parameters ASR PSNR $\bar{n}_{it}$ Time(s)
JMA (0.05,200) 0.95 24.08 48.61 196.55
(0.1,200) 0.96 23.87 31.60 136.07
(0.2,200) 1.00 23.40 16.92 62.36
(0.5,200) 1.00 22.34 12.04 45.53
(1,200) 1.00 20.32 8.69 33.58
C&W (1e-3,10,100) 0.01 26.64 53.00 48.96
(1e-3,10,200) 0.11 22.17 93.82 85.91
(1e-3,10,500) 0.37 20.78 176.81 163.32
(1e-3,10,1000) 0.42 20.87 328.05 327.64
(1e-1,10,2000) 0.52 20.59 910.73 841.25
ECOC attack (1e-3,10,100) 0.18 21.11 92.28 131.46
(1e-3,10,200) 0.38 21.32 181.63 259.54
(1e-3,10,500) 0.59 21.58 488.73 697.33
(1e-3,10,1000) 0.69 21.87 967.56 1387.41
(1e-1,10,2000) 0.73 22.24 1921.29 2754.74
ML-CW (0.01, 10, 1000) 0.64 18.47 297.53 1679.72
LOTS $n_{it,max}=2000$ 0.78 23.66 242.23 600.14
MLA-LP $n_{it,max}=1000$ 0.00 NA NA NA (1079 s)
MLA-LP $n_{it,max}=3000$ 0.00 NA NA NA ($>$3600 s)
6.3 Multi-label classification
Table VI. shows the results we got in the case of multi-label classification when the target label vector is chosen randomly among those contained in the training dataset.
The average number of labels targeted by the attacks is 2.1999Most of the images in the VOC2012 dataset contains one or very few labeled objects, hence the label vectors have few 1’s and choosing the target label from the training dataset results in few bit changes. We see that both JMA and ML-C&W attack get ASR = 1. However, JMA has an advantage over ML-C&W both in terms of PSNR (around 3dB) and, above all, in terms of computational complexity, being approximately 64 times faster than ML-C&W. In contrast to what one would expect, LOTS achieves much worse performance, with ASR = 0.66 and a lower PSNR. However, the average bASR is $0.979$, that confirms that the method has some effectiveness. We believe that the reason for the not so good results of LOTS relies on the construction of the target vector performed by the method. In fact, compared to multi-class classification where images belonging to the same class have similar content, e.g. images showing the same traffic sign, in the multi-label classification case, images sharing the same label vector may be very different from each other, and the differences between them grow with the amount of categories included in the label vector. For instance, an image of a crowded city street and an image with footballers on a soccer field are both instances of the people category, with the ’people’ label equal to 1. As LOTS chooses the target output vector by averaging the logits of the target images, the resulting target vector might not be plausible when the logits of the target images are significantly different. Finally, the performance of MLA-LP are also poor, with a much lower ASR w.r.t. JMA and ML-C&W. Moreover, the computational cost of the attack is very high. These results are not in contrast with the results reported in [11] for the same dataset, given that in [11] the method is validated for target attacks that change only one bit of the original label vector, while the average number of bits changed in our experiments is larger than 2.
Tables VII through IX show the results in a challenging scenario where the target vector labels are not chosen from the training/test set but are obtained by randomly changing a certain number of bits of the label vector of the to-be-attacked image. Specifically, Tables VII, VIII and IX correspond to attacks inducing respectively 5, 10 and 20 bit errors (in the latter case all the bits should be changed). The results are reported for JMA, ML-C&W and MLA-LP. As we said LOTS can not be applied in this case, since this method requires the availability of a number of samples corresponding to the target label vector - that, in principle, in this case may not even correspond to a plausible image.
Looking at the results in Table VII, we see that in the 5 bit case JMA can achieve ASR = 0.95 (bASR = 0.99), which is 5% higher than ML-CW with an average attack time with is approximately 6 times lower, while MLA-LP is totally ineffective (ASR = 0).
For the case of 10 bit and 20 bit errors, we also report the results with larger $n_{it,max}$ for both JMA and ML-C&W. In fact, achieving a successful attack in the case where half of the bits and, even worse, all the bits are changed is extremely hard, and more iterations are necessary. From Table VIII we see that, in the 10 bit case, JMA can still achieve an ASR close to 0.90, at the price of a longer time necessary to run the attack, with an average number of iterations $\bar{n}_{it}$ that goes above 240. The C&W attack achieves a lower ASR with a bit larger PSNR when $n_{it,max}$ is raised to 5000, with a very large computational cost. Fig. 3 shows some examples of images successfully attacked by JMA in the 10-bit error case, when half of the labels are modified.
As shown in Table IX, the ASR drops in the 20 bit case, namely, for an attack that tries to modify all the 20 components of the label vector. Arguably, this represents a limit and extremely challenging scenario. We see that, by increasing $n_{it,max}$ to 2000, the JMA attack is successful 26% of the times, at the price of a high computational cost, while ML-C&W is never successful. Moreover, it is interesting to observe that JMA can achieve a bASR of 0.68, meaning that almost 70% of the labels are modified by the attack, while the bASR for ML-C&W is below 0.1. Not surprisingly, the PSNR is low in this case.
Figure 3: Examples of JMA attacked images for which the attack can successfully flip 10 targeted bits of the label vector in the multi-label (VOC2012) classification task. For the car: the original image has label vector ( ’car’=1 , ‘person’ =0, ‘cat’ = 0, ‘cow’ = 0, ‘dog’ = 0, ‘sheep’ = 0, ‘bike’ = 0, ‘chair’ = 0, ‘potted plant’ = 0, ‘monitor’ = 0, ’aeroplane’ = 0, ‘bird’ = 0, ‘horse’ =0, ‘boat’ = 0, ‘dining table’ = 0, ’train’ = 0,…..), while the JMA attacked image is classified as ( ’car’=0 , ‘person’ =1, ‘cat’ = 1, ‘cow’ = 1, ‘dog’ = 1, ‘sheep’ = 1, ‘bike’ = 1, ‘chair’ = 1, ‘potted plant’ = 1, ‘monitor’ = 1, ’aeroplane’ = 0, ‘bird’ = 0, ‘horse’ =0, ‘boat’ = 0, ‘dining table’ = 0, ’train’ = 0,…..). For the airplane: the original image has label vector ( ’car’=0 , ‘person’ =0, ‘cat’ = 0, ‘cow’ = 0, ‘dog’ = 0, ‘sheep’ = 0, ‘bike’ = 0, ‘chair’ = 0, ‘potted plant’ = 0, ‘monitor’ = 0, ’aeroplane’ = 1, ‘bird’ = 0, ‘horse’ =0, ‘boat’ = 0, ‘dining table’ = 0, ’train’ = 0,…..), while the label vector of the JMA attacked image is ( ’car’=0 , ‘person’ =1, ‘cat’ = 1, ‘cow’ = 1, ‘dog’ = 0, ‘sheep’ = 0, ‘bike’ = 0, ‘chair’ = 0, ‘potted plant’ = 1, ‘monitor’ = 1, ’aeroplane’ = 0, ‘bird’ = 0, ‘horse’ =1, ‘boat’ = 1, ‘dining table’ = 1, ’train’ = 0,…..). The corresponding PSNR value is also reported.
TABLE VI: Results of the attacks against multi-label classification on the VOC2012 dataset. On the average the attacks are asked to flip 2.1 labels.
Parameters ASR bASR PSNR $\bar{n}_{it}$ Time(s)
JMA (0.05,200) 1.00 1.0000 56.54 29.94 25.97
(0.1,200) 1.00 1.0000 55.78 17.50 15.39
(0.5,200) 1.00 1.0000 49.66 7.51 7.32
(1,200) 1.00 1.0000 44.86 6.42 6.45
ML-C&W (0.01, 10, 1000) 1.00 1.0000 52.03 180.93 469.90
LOTS 2000 0.66 0.9787 50.91 437.20 268.16
MLA-LP 100 0.58 0.9560 51.17 50.44 1361.29
TABLE VII: Results for VOC2012 classification with random choice of the target label vector (5 bits out of 20 are flipped).
Parameters ASR bASR PSNR $\bar{n}_{it}$ Time(s)
JMA (0.05,200) 0.49 0.9032 49.18 129.85 139.31
(0.1,200) 0.82 0.9687 47.48 110.21 106.98
(0.2,200) 0.95 0.9942 45.26 87.96 79.05
ML-C&W (0.01, 10, 1000) 0.90 0.9883 45.29 113.35 420.21
MLA-LP 100 0 0.4264 NA NA NA
TABLE VIII: Results for VOC2012 classification with random choice of the target label vector (10 bits out of 20 are flipped).
Parameters ASR bASR PSNR $\bar{n}_{it}$ Time(s)
JMA (0.05,200) 0.01 0.5714 48.16 161.00 157.41
(0.1,200) 0.06 0.6700 45.62 163.84 156.04
(0.2,200) 0.23 0.8112 42.77 163.40 162.63
(0.5,200) 0.31 0.8842 35.60 161.36 158.88
(0.1, 500) 0.56 0.8892 43.61 312.42 397.42
(0.5, 500) 0.88 0.9805 35.66 245.24 278.02
ML-C&W (0.01, 10, 1000) 0.295 0.8697 45.09 270.17 498.63
(0.01, 10, 2000) 0.4186 0.9156 44.46 451.78 939.53
(0.1, 10, 2000) 0.475 0.9147 44.73 433.75 934.58
(0.5, 10, 2000) 0.475 0.9147 44.78 430.96 908.92
(0.1, 10, 3000) 0.575 0.9331 44.81 581.53 1461.05
(0.1, 10, 5000) 0.72 0.9557 44.74 775.97 2183.79
MLA-LP 100 0 0.5167 NA NA NA
TABLE IX: Results for VOC2012 classification with random choice of the target label vector. Case of 20 bit errors (all labels are changed by the attack).
Parameters ASR bASR PSNR $\bar{n}_{it}$ Time(s)
JMA (0.05,200) 0 0.1109 NA NA NA
(0.1,200) 0 0.1276 NA NA NA
(0.2,200) 0 0.1475 NA NA NA
(0.5,200) 0 0.1845 NA NA NA
(0.5,500) 0 0.3288 NA NA NA
(0.5,1000) 0.03 0.5068 30.86 938.16 980.82
(0.5,2000) 0.26 0.6795 29.43 1484.34 1677.49
ML-C&W (0.01, 10, 1000) 0 0.0659 NA NA NA
(0.5, 10, 2000) 0 0.0882 NA NA NA
7 Discussion
Our experiments showed that JMA requires a small number of iterations to find an adversarial example, much smaller than the other algorithms, and gets an ASR that is generally higher (in some cases, much higher) for a similar distortion. It is worth stressing that the number of updates $\bar{n}_{it}$ reported in the tables takes into account also the steps of the binary search carried out at the end of the iterations (see Section 5.4). Specifically, $it=(v-1)+n_{bs}$, where $v$ is the number of ’for’ loops, that is, the number of image updates (see Algorithm 1), and $n_{bs}$ is set equal to $6$. Then, the $\bar{n}_{it}$ values reported in the various tables reveal that JMA can indeed find an adversarial image in one shot in many cases, see for instance Table IV for CIFAR-10 ECOC classification and Table VI for the multi-label classification where $\bar{n}_{it}$ is lower than 7 for the attack settings with larger $\epsilon$ [^10] Note that, even when a one-shot attack is possible, a perturbation applied to the image with a weak strength $\epsilon$ may not result in an adversarial image, thus requiring that more iterations are run.. This confirms a-posteriori the validity of one of the most basic assumptions underlying our approach, namely the local linear behavior, that, when it holds, allows to perform the attack in one-shot. When the linear approximation does not hold in the close vicinity of the initial point, and possibly of other points obtained after the updating, the perturbation is added with a small $\epsilon$ in order to change the initial point, and the Jacobian is recomputed (see Algorithm 1). By inspecting the tables (see for instance Table II), we can see that, when this happens, it is often preferable to use a not-too-small $\epsilon$ so that the point is moved farther from the initial point, where the behavior of the local function can be closer to linear, instead of remaining in the vicinity of the initial point where the local approximation may still be not satisfied. Doing so, a high ASR can be achieved with a reduced number of iterations and a similar PSNR.
8 Conclusions
We have proposed a general, nearly optimal, targeted attack, that can solve the original formulation of the adversarial attack by Szegedy et al. under a first order approximation of the network function. The method resorts to the minimization of a Mahalanobis distance term, which depends on the Jacobian matrix taking into account the effort necessary to move the output representation of the image in a given direction. By exploiting the Wolfe duality, the minimization problem is reduced to a non-negative least square (NNLS) problem, that can be solved numerically. The experiments show that the JMA attack is effective against a wide variety of DNNs adopting different output encoding schemes, including networks using error correction output encoding (ECOC) and multi-label classification, outperforming existing attacks in terms of larger attack success rate, lower distortion and lower complexity. In particular, in the multi-label case, our experiments show that JMA is capable to simultaneously change up to 10 (random) labels of the output label vector for VOC2012 classification with a very high ASR (and, in some cases, it can also change all the 20 output labels), a capability that is completely out of reach of the attacks proposed so far. JMA remains effective also in the case of one-hot encoding, where it has much reduced computational complexity with respect to the C&W attack.
Future work will focus on exploiting the JMA attack in a black-box attack scenario, to develop powerful targeted attacks with certain transferability properties. Exploiting the proposed attack for adversarial training, to increase robustness of general classification models against targeted attacks, is also another interesting research direction.
Acknowledgment
This work has been partially supported by the Italian Ministry of University and Research under contract PRIN 2017 2017Z595XS-001 (PREMIER project).

### v2

@@
 Most of the approaches proposed so far to craft targeted adversarial examples against Deep Learning classifiers are highly suboptimal and typically rely on increasing the likelihood of the target class, thus implicitly focusing on one-hot encoding settings.
-In this paper, we propose a more general, theoretically sound, targeted attack that resorts to the minimization of a Jacobian-induced MAhalanobis distance (JMA) term, taking into account the effort (in the input space) required to move the latent space representation of the input sample in a given direction.
+In this paper, a more general, theoretically sound, targeted attack is proposed, which resorts to the minimization of a Jacobian-induced Mahalanobis distance term, taking into account the effort (in the input space) required to move the latent space representation of the input sample in a given direction.
 The minimization is solved by exploiting the Wolfe duality theorem, reducing the problem to the solution of a Non-Negative Least Square (NNLS) problem.
-The proposed algorithm provides an optimal solution to a linearized version of the adversarial example problem originally introduced by Szegedy et al.
-[1].
-The experiments we carried out confirm the generality of the proposed attack which is proven to be effective under a wide variety of output encoding schemes.
-Noticeably, the JMA attack is also effective in a multi-label classification scenario, being capable to induce a targeted modification of up to half the labels in a complex multilabel classification scenario with 20 labels, a capability that is out of reach of all the attacks proposed so far.
-As a further advantage, the JMA attack usually requires very few iterations, thus resulting more efficient than existing methods.
+The proposed algorithm (referred to as JMA) provides an optimal solution to a linearised version of the adversarial example problem originally introduced by Szegedy et al.
+The results of the experiments confirm the generality of the proposed attack which is proven to be effective under a wide variety of output encoding schemes.
+Noticeably, JMA is also effective in a multi-label classification scenario, being capable to induce a targeted modification of up to half the labels in complex multi-label classification scenarios, a capability that is out of reach of all the attacks proposed so far.
+As a further advantage, JMA requires very few iterations, thus resulting more efficient than existing methods.
 Index Terms:
 Adversarial Examples, Deep Learning Security, Adversarial Machine Learning, Multi-Label Classification, Mahalanobis Distance, Non-Negative Least Square Problems
-publicationid: pubid: $0$–$00$ 2021 IEEE
-1 Introduction
+I Introduction
 Adversarial examples, namely, quasi-imperceptible perturbations capable to induce an incorrect decision, are a serious threat to deep-learning classifiers [1, 2].
 Since the publication of the seminal work by Szegedy et al.
-[1] in which the existence of adversarial examples was first observed, a large number of gradient-based methods have been proposed to implement adversarial attacks against Deep Neural Networks (DNNs) in white-box and black-box scenarios.
-Most of the attacks proposed so far rely on greedy algorithms that permit to find an effective adversarial example in a reasonable amount of time [2, 3, 4, 5, 6, 7].
+[1] in which the existence of adversarial examples was first observed, a large number of gradient-based methods have been proposed to implement adversarial attacks against Deep Neural Networks (DNNs) in white-box and black-box scenarios [2, 3, 4, 5].
 Most attacks create the adversarial examples by minimizing a cost function subject to a constraint on the maximum perturbation introduced in the image.
-Moreover, they focus on classifiers based on one-hot encoding.
+Moreover, they focus on single-label classifiers based on one-hot encoding.
 In this setting, the final activation layer consists in the application of an activation function (usually a softmax) and a normalization, mapping the last layer outputs, called logits, into a probability vector, associating a probability value to each class.
@@
 As a consequence, in the targeted case, adversarial examples focus on raising the probability value of the target class.
-This is obviously the best strategy with one-hot encoding in the presence of a softmax function, however, this strategy is no more optimal in the presence of output encoding schemes based on channel coding, and in the case of multi-label classification.
+This is obviously the best strategy with single-label classification however, this strategy is not optimal in general, e.g., in the presence of output encoding schemes based on channel coding, and in the case of multi-label classification.
 A more flexible attack working at the logits level has been proposed by Carlini and Wagner [6].
@@
 In this case, all the logits have to be modified simultaneously.
-Crafting adversarial examples for networks adopting a generic output encoding scheme, and for multi-label classification, is a very challenging task, that, to the best of our knowledge, has not been given much attention, with the exception of some scattered works proposing suboptimum approaches [10, 11].
+Crafting adversarial examples for networks adopting a generic output encoding scheme, and for multi-label classification, is a very challenging task, that, to the best of the authors’ knowledge, has not been given much attention, with the exception of some scattered works proposing suboptimum approaches [9, 10], [11].
 From the attacker side, the difficulties of applying an optimal attack in this case are due to the fact that the labels are not mutually exclusive (like in the one- hot case), and decreasing or increasing the network output in correspondence of some nodes may change the values taken by the other nodes in an unpredictable way.
 To overcome the above drawbacks, we introduce a new, theoretically sound, targeted adversarial attack, named Jacobian-induced Mahalanobis distance Attack (JMA).
-JMA solves the original formulation of the adversarial example problem introduced by Szegedy et al.
-[1], which aims at minimizing the strength of the adversarial perturbation subject to the constraint that the image is classified as belonging to the target class.
+JMA solves the original formulation of the adversarial example problem introduced in [1], which aims at minimizing the strength of the adversarial perturbation subject to the constraint that the image is classified as belonging to the target class.
 The solution is found via a two-step procedure: i) given a target point in the output space, the optimum perturbation moving the input sample to the target point is first determined, under a linear assumption on the effect of the distortion on the model output; ii) then, the target point that minimizes the perturbation introduced by the attack is determined by minimizing the Mahalanobis distance induced by the Jacobian matrix of the network input-output function.
-By exploiting the Wolfe duality theorem, we reduce the problem to the solution of a Non-Negative Least Square (NNLS) problem, that can be efficiently solved numerically.
+By exploiting the Wolfe duality theorem, the problem is reduced to the solution of a Non-Negative Least Square (NNLS) problem, that can be efficiently solved numerically.
 The optimal perturbation derived in this way is then applied to the input image.
-Ideally, the JMA attack should produce the desired adversarial image in just one-shot.
-However, due to the local optimality of the proposed approach, in practice, it is sometimes necessary to carry out some iterations to obtain a valid adversarial example, recomputing the Jacobian matrix every time.
-The experiments we carried out confirm that the proposed attack is more efficient than state-of-the-art attacks from a computational perspective, requiring a much lower number of iterations to carry out the attack.
+Ideally, JMA should produce the adversarial image in just one-shot.
+However, due to local optimality, in practice, it is sometimes necessary to carry out some iterations to obtain a valid adversarial example, recomputing the Jacobian matrix every time.
+The results of the experiments confirm that the proposed attack is more efficient than state-of-the-art attacks from a computational perspective, requiring lower iterations to carry out the attack.
 State-of-the-art attacks are also outperformed in terms of distortion and attack success rate.
-The effectiveness of the proposed adversarial attack method is particularly evident when it is used to attack networks adopting Error Correction Output Coding (ECOC) [9] and in multi-label classification scenarios, where the JMA attack is capable to simultaneously change in a desired way up to 10 out of 20 labels of the output label vector, a capability which is out of reach of the algorithms proposed so far.
-We also verified that the method has a very good behavior in the one-hot encoding scenario, where it achieves performance that are comparable to the state-of-the-art with a significant reduction of the computing time.
-In summary, the main contributions of the paper are:
-  • We introduce a new targeted adversarial attack, named Jacobian-induced Mahalanobis distance Attack (JMA).
-The attack resorts to the minimization of a Jacobian-induced MAhalanobis distance (JMA) term, with the Jacobian matrix taking into account the effort (in the input space) required to move the latent space representation of the input sample in a given direction.
+The effectiveness of the proposed adversarial attack method is particularly evident when it is used to attack networks adopting ECOC and in particular in multi-label classification scenarios, where JMA is capable to simultaneously change in a desired way up to 10 out of 20 labels of the output label vector, a capability which is out of reach of the algorithms proposed so far.
+We also verified the very good behavior of JMA in the one-hot encoding scenario, where it achieves performance that are comparable to the state-of-the-art with a significant reduction of the computing time.
+In summary, the main contributions of this paper are:
+  • We propose a new targeted adversarial attack for general classification frameworks, named Jacobian-induced Mahalanobis distance Attack (JMA).
+The attack resorts to the minimization of a Jacobian-induced Mahalanobis distance term, with the Jacobian matrix taking into account the effort (input space) to move the latent space representation of the input sample in a given direction.
 In theory, the algorithm is one-shot.
   • We solve the constrained minimization of the Mahalanobis distance by exploiting the Wolfe duality theorem, reducing the problem to the solution of a non-negative least square (NNLS) problem, that can be solved numerically.
-  • We validate the new attack considering several datasets (CIFAR-10, GTSRB, MNIST, VOC2012) and networks adopting different encoding schemes, namely the ECOC scheme, multi-label classification, and also one-hot encoding.
-Our experiments confirm that the proposed attack is very general and can work in all these cases, being very efficient, often requiring only very few iterations to attack the image.
-  • We compare JMA against several state-of-the-art methods, showing that the proposed attack is more efficient and requires a much lower number of iterations.
-In the cases of ECOC and multi-label classification, it also achieves lower distortion and higher attack success rate.
-In particular, in the multi-label classification scenario, the proposed attack is capable to change up to half labels of the sample output vector, even when the number of labels increases.
-The rest of the paper is organized as follows: in Section 2, we provide the main formalism and describe the various classification frameworks addressed in the paper.
-In Section 3, we review the most relevant adversarial attacks introduced so far, with particular attention to the attacks against DNNs adopting with output channel encoding.
-Then, in Section 4, we present the JMA adversarial attack.
-The experimental methodology and setting are described in Section 5.
-Sections 6 and 7 report and discuss the results of the experiments.
-The papers ends in Section 8 with our final conclusions.
-2 Background and Notation
+  • We validate the new attack on various datasets (CIFAR-10, GTSRB, MNIST, VOC2012, MS-COCO, NUS-WIDE, and ImageNet) and networks adopting different encoding schemes, namely ECOC, multi-label, and one-hot encoding.
+Validation involves different types of architectures, including CNNs, Transformers and large visual models.
+The experiments confirm that the proposed attack is very general and can work in all these cases, being very efficient, often requiring only very few iterations to attack an image.
+  • We compare our new approach against several state-of-the-art attacks, showing that JMA is more efficient and requires less iterations.
+In particular, in the multi-label classification scenario, JMA is capable to change up to half labels of the output vector, even when the number of labels increases.
+The rest of the paper is organized as follows: Section II gives an overview of the related work.
+Section III provides the main formalism and introduces the various classification frameworks.
+In Section IV, the most relevant adversarial attacks that can be applied against DNNs adopting output encoding are reviewed.
+Then, in Section V, the details of the JMA adversarial attack are described.
+The experimental methodology and setting are described in Section VI, while section VII reports and discusses the results of the experiments.
+The papers ends in Section VIII with some final remarks.
+II Related works
+Adversarial attacks can be categorized in two main groups, namely white-box and black-box methods [5].
+Starting from Szegedy et al.
+[1] work, research on white-box attacks has mainly focused on the development of gradient-based approaches that can reduce the complexity of the attack.
+Many greedy algorithms have been proposed that permit to find an effective adversarial example in a reasonable amount of time [2, 3, 12, 13, 14, 6].
+The Fast Sign Gradient Method (FGSM) method [2] obtains an adversarial perturbation in a computationally efficient way by considering the sign of the gradient of the output with respect to the input image.
+In [12], an iterative version of FGSM is introduced by applying FGSM multiple times, with a smaller step size, each time by recomputing the gradient.
+This method is often referred to as I-FGSM, or Basic Iterative Method (BIM).
+Another attack similar to FGSM is the projected gradient descent (PGD) attack [13], that can be regarded to as a multi-step extension of the FGSM attack where the clip operation performed by BIM on the gradient (to force the solution to stay in the [0,1] range) is replaced by gradient projection.
+Carlini and Wagner (C&W) [6] propose a more flexible method, working at the logits level, which can mitigate the gradient vanishing problem and improve performance in many cases.
+Recently, the AutoPGD and AutoAttack [15] attacks have also been proposed.
+Such attacks improve PGD by designing an approach that automatically chooses the most suitable step size and perturbation size (in the case of AutoAttack) at every attack iteration, based on the behaviour of the objective function, and by using a different loss function, which improves performance and reduces the gradient vanishing problem.
+Among the other white-box gradient-based attacks it is worth mentioning the Jacobian-based Saliency Map Attack (JSMA) [3] and the DeepFool attack [14].
+JSMA [3] consists of a greedy iterative procedure which relies on forward propagation to compute, at each iteration, a saliency map, indicating the pixels that contribute most to the classification, while DeepFool [14] is an efficient iterative attack that considers the minimal perturbation with respect to a linearized classifier, stopping the attack when the boundary is crossed.
+A large segment of recent literature has also focused on the development of attacks that can work in real-world settings where the exact victim model is unknown.
+This includes transfer-based attacks, e.g.
+[4, 16, 17], which retain part of their effectiveness even against DNN models other than the one targeted by the attack, as well as black-box attacks operating under the assumption that the victim model can be queried a limited number of times [18, 19].
+All the above methods focus on single-label classifiers adopting one-hot encoding schemes.
+The problem of crafting adversarial attacks against classifiers adopting different output encoding schemes is a challenging one that, to the best of the authors’ knowledge, has not been much studied, with the exception of a few scattered works.
+Song et.
+al [9] first addressed this problem and extended the C&W and the DeepFool attacks to a multi-label setting.
+These extended algorithms are referred as ML-C&W and ML-DF.
+Inspired by ML-DF, another approach to implement an adversarial attack in the multi-label case, named Multi-Label Attack via Linear programming (MLA-LP), has been proposed in [10].
+Recently, [11] proposed a fundamentally different multi-label attack method that enforces semantic consistency across all predicted labels in the adversarial image.
+The approach introduced in [11] leverages an efficient search algorithm over a knowledge graph that captures label dependencies.
+Beyond multi-label classification, a white-box attack tailored against DNNs adopting the ECOC encoding scheme has been proposed in [20].
+A completely different approach, that can be applied to any network regardless of the encoding scheme, often leading to a large attack distortion in the input space, is the Layerwise Origin-Target Synthesis (LOTS) attack [21].
+Due to their relevance for our research, the attacks targeting classifiers adopting output encoding are presented in detail in Section IV-B.
+III Background and Notation
 In this section, we review the main output encoding schemes for DNN classifiers.
 Let $x\in\mathbb{R}^{m}$ denote the input of the network, and $y(x)$ (or simply $y$) the class $x$ belongs to.
-We indicate with $l$ the number of classes.
+Le $l$ be the number of classes.
 To classify $x$, the neural network first maps $x$ into a reduced dimensional space $\mathbb{R}^{n}$ ($n<m$).
@@
 Classification in favour of one of the $l$ classes is obtained by applying a function $\phi$ to $f(x)$, the exact form of $\phi$ depending on the output encoding scheme used by the network.
-In the following, we denote by $z=(z_{1},z_{2},...,z_{n})^{T}$ the vector with the logit values, that is, the values of the network nodes before the final activation function, which is responsible to map the output of the penultimate layer of the network into the [0,1] (sometimes [-1, 1]) range.
-Given an image $x$, the notation $f^{-j}(x)$ is used to refer to the internal model representation at layer $L-j$ ($L$ being the total number of layers of the network).
+In the following, $z=(z_{1},z_{2},...,z_{n})^{T}$ denotes the vector with the logit values, that is, the values of the network nodes before the final activation function, which is responsible to map the output of the penultimate layer of the network into the [0,1] (sometimes [-1, 1]) range.
+Given an image $x$, the notation $f^{-j}(x)$ is used to refer to the internal model representation at layer LjL-j ($L$ being the total number of layers of the network).
 With this notation, $z(x)=f^{-1}(x)$.
-2.1 One-hot encoding
+III-A One-hot encoding
 In the case of networks adopting the one-hot encoding scheme, the number of output nodes corresponds to the number of classes ($l=n$), and $c_{k}$ is a binary vector with all 0’s except for position $k$, where it takes value 1.
@@
 for $k=1,..,l$. This allows to interpret $f_{k}(x)$ as the probability assigned to class $k$, and the final prediction is made by letting $\phi(x)={\arg\max}_{k}\hskip 2.84544ptf_{k}(x)$.
-Typically, in the one-hot encoding case, training is carried out by minimizing the categorical cross-entropy loss defined as $\mathcal{L}(x,y)=-\sum_{i=1}^{l}c_{yi}\log(f_{i}(x))=-\log f_{y}(x)$ [^1] We remind that $y$ denotes the class sample $x$ belongs to..
-2.2 Error-correction-output-coding (ECOC)
+Typically, in the one-hot encoding case, training is carried out by minimizing the categorical cross-entropy loss defined as $\mathcal{L}(x,y)=-\sum_{i=1}^{l}c_{yi}\log(f_{i}(x))=-\log(f_{y}(x))$.
+III-B Error-correction-output-coding (ECOC)
 Sometimes the output classes are encoded by using the codewords of a channel code.
@@
 Each element of $C$ can take values in $\{-1,1\}$.
-To compute the output of the network the logits are mapped into the $[-1,1]$ range by means of an activation function $\sigma()$, that is $f(x)=\sigma(z)$, where $\sigma()$ is applied element-wise to the logits.
+To compute the output of the network the logits are mapped into the [1,1][-1,1] range by means of an activation function $\sigma()$, that is $f(x)=\sigma(z)$, where $\sigma()$ is applied element-wise to the logits.
 A common choice for $\sigma()$ is the $\tanh$ function.
@@
 In the ECOC case, training is usually carried out by minimizing the hinge loss function, defined as $\mathcal{L}(x,y)=\sum_{i=1}^{n}\max(1-c_{yi}f_{i}(x),0)$.
-2.3 Multi-label classification
-In this paper, we refer to multi-label classification as a scenario wherein the classifier is asked to decide about the presence of absence within the image of $n$ image characteristics or features.
+III-C Multi-label classification
+Multi-label classification is the scenario wherein the classifier is asked to decide about the presence or absence within the image of $n$ image characteristics or features.
 For example, the classifier may be asked to detect the presence of $n$ possible classes of objects, or decide about $n$ binary attributes (like indoor/outdoor, night/day, sunny/rainy).
@@
 In such a case, $C$ contains only the allowed combinations, somewhat playing the same role of channel coding.
-In this paper, by multi-label classification, we indicate a situation where all combinations of labels are possible.
+In this paper, multi-label classification indicates a situation where all combinations of labels are possible.
 In the case of multi-label classification, given the logit vector $z$, the activation function $\sigma()$ is applied element-wise to the components of $z$, and the prediction on each label component is made component-wise.
-Assuming that $c_{k}\in\{0,1\}^{n}$ for any $k$, and that the logistic function is applied to the logits, decoding corresponds to 0.5-thresholding each output score independently [^2] If $c_{k}\in\{-1,1\}^{n}$, the tanh activation is applied followed by 0-thresholding..
-For the loss function, a common choice is the binary cross-entropy loss $\mathcal{L}(x,y)=-\sum_{i=1}^{n}(c_{yi}\log(f_{i}(x))+(1-c_{yi})\log(1-f_{i}(x)))$.
-3 Adversarial attacks against DNNs
+Assuming that $c_{k}\in\{0,1\}^{n}$ for any $k$, and that the logistic function is applied to the logits, decoding corresponds to 0.5-thresholding each output score independently [^1] If $c_{k}\in\{-1,1\}^{n}$, tanh activation followed by 0-thresholding is applied..
+For the loss function, a common choice is the (multi-label) binary cross-entropy loss $\mathcal{L}(x,y)=-\sum_{i=1}^{n}(c_{yi}\log(f_{i}(x))+(1-c_{yi})\log(1-f_{i}(x)))$.
+IV Adversarial attacks against DNNs
 The vulnerability of deep neural networks (DNNs) to adversarial samples has been first pointed out by Szegedy et al.
 in [1].
-When the goal of the attacker is to induce a generic misclassification, we talk about an untargeted attack.
-When the misclassification is in favor of a specific class, we say that the attack is a targeted one.
+When the goal of the attacker is to induce a generic misclassification, the attack is referred to as untargeted Conversely, if the misclassification aims at a specific class, the attack is referred to as targeted.
 According to [1], the generation of a targeted adversarial example can be formally described as
-$&\textrm{minimize}\quad||\delta||_{2}\\ &\textrm{s.t.}\quad\phi(x+\delta)=t\neq\phi(x)\\ &\textrm{and}\quad x+\delta\in[0,1]^{m}\end{aligned},$ (3)
-where $||\delta||_{2}$ denotes the $l_{2}$ norm of the perturbation $\delta$, and $t$ is the target class of the attack.
+[…]
+where $\delta||_{2}$ denotes the $l_{2}$ norm of the perturbation $\delta$, and $t$ is the target class of the attack.
 The same formulation holds for the untargeted case, with the first constraint replaced by $\phi(x+\delta)\neq\phi(x)$.
 The constraint $x+\delta\in[0,1]^{m}$ makes sure that the resulting adversarial example is a valid input.
-3.1 Basic adversarial attacks
-Solving the minimization in (3) is generally hard, then the adversarial examples are determined by solving the simplified problem described below:
-$\min_{\delta}$ $\hskip 5.69046pt\mathcal{L}({x+\delta},t)+\lambda\cdot\left\|\delta\right\|_{2}^{2}$ (4)
-s.t. $\hskip 5.69046ptx+\delta\in[0,1]^{m}$
-where $\mathcal{L}({x+\delta},t)$ is the loss function under the target class $t$, and $\lambda$ is a parameter balancing the importance of the two terms.
-The problem in (4) can be solved by the box-constraint L-BFGS method [1].
-Specifically, a line search is carried out to find the value of $\lambda>0$ for which the solution of (4) satisfies the adversarial condition, that is $\phi(x+\delta)=t$.
-Usually, the categorical cross entropy loss is considered.
-From (4), it is immediate to see that, in the one-hot case, when the categorical cross entropy loss is considered, only the target class contributes to the loss term, and then the attack only cares about increasing $f_{t}(x)$, regardless of the other scores, all the more that, due to the presence of the softmax, this also implies decreasing the other outputs.
-To reduce the complexity of the L-BFGS attack, several suboptimal solutions have been proposed, that consider the problem - equivalent to (4) - of minimizing the loss function $\mathcal{L}$ subject to a constraint on the perturbation $\delta$.
-The Fast Sign Gradient Method (FGSM) method [2] obtains an adversarial perturbation in a computationally efficient way by considering the sign of the gradient of the output with respect to the input image.
-The FGSM attack is a one-step attack whose strength can be modulated only by increasing the value of the term multiplying the gradient sign.
-In [4], an iterative version of FGSM is introduced by applying FGSM multiple times, with a smaller step size, each time by recomputing the gradient.
-This method is often referred to as I-FGSM, or Basic Iterative Method (BIM).
-Another attack similar to the FGSM attack is the projected gradient descent (PGD) attack [5], that can be regarded to as a multi-step extension of the FGSM attack where the clip operation performed by BIM on the gradient (to force the solution to stay in the [0,1] range) is replaced by the gradient projection.
-Other gradient-based attacks are the Jacobian-based Saliency Map Attack (JSMA) [3] and the DeepFool attack [6].
-JSMA [3] consists of a greedy iterative procedure which relies on forward propagation to compute, at each iteration, a saliency map, indicating the pixels that contribute most to the classification.
-In such attack, the Jacobian matrix of the network function is used to compute the forward derivative and build the saliency map.
-A variant of this attack, called Maximal JSMA, has been proposed in [15].
-DeepFool [6] is an efficient iterative attack that considers the minimal perturbation with respect to a linearized classifier.
-In this way, the current version of the adversarial example is projected onto the separating affine hyperplane at each iteration, until the classification boundary is crossed.
-The above gradient-based methods often fail due to the gradient vanishing problem [7].
-A solution to overcome this issue has been proposed in [16], where the values of the logits are scaled by a temperature parameter.
-However, this approach lacks flexibility and the suitable distillation temperature parameter strongly depends on the network.
-A more flexible and efficient method working at the logits level has been proposed by Carlini and Wagner (C&W) [7].
-In the C&W attack the loss term is replaced by a properly designed function depending on the difference between the logit of the target class and the largest among the logits of the other classes (since the logits do not sum one, two logits at a time must be considered).
+IV-A Basic adversarial attacks
+Solving the minimization in (3) is generally hard, then the adversarial examples are determined by solving the simplified problem where a functional $\mathcal{L}({x+\delta},t)+\lambda\cdot\left\|\delta\right\|_{2}^{2}$ is minimized, where $\mathcal{L}({x+\delta},t)$ is the loss function (usually, the categorical cross entropy loss) under the target class $t$, and $\lambda$ is a parameter balancing the two terms.
+This problem can be solved by the box-constraint L-BFGS method [1].
+Specifically, a line search is carried out to find the value of $\lambda>0$ for which the solution satisfies the adversarial condition, that is $\phi(x+\delta)=t$.
+It is immediate to see that, in the one-hot case, when the categorical cross entropy loss is considered, only the target class contributes to the loss term, and then the attack only cares about increasing $f_{t}(x)$, regardless of the other scores, all the more that, due to the presence of the softmax, this also implies decreasing the other outputs.
+To reduce the complexity of the L-BFGS attack, several suboptimal solutions have been proposed, considering the problem of minimizing the loss function $\mathcal{L}$ subject to a constraint on the perturbation $\delta$, like I-FGSM (or BIM) [12] and PGD [13].
+A flexible and efficient method working at the logits level has been proposed by Carlini and Wagner [6].
+In the C&W attack in fact, the loss term is replaced by a properly designed function depending on the difference between the logit of the target class and the largest among the logits of the other classes.
 The perturbation is kept in the valid range by properly modifying the objective function.
-In this way, the following unconstrained problem is considered and solved (for the case of $L_{2}$ distortion metric):
-$\min_{w}$ $\hskip 5.69046pt||\delta(w)||_{2}^{2}+$
+In this way, the following unconstrained problem is considered and solved (for the case of $L_{2}$ metric):
+[…]
 $\lambda\cdot\max(\max_{i\neq t}z_{i}(x+\delta(w))-z_{t}(x+\delta(w)),-\kappa),$ (4)
@@
 Since logits are more sensitive to modifications of the input than the probability values obtained after the softmax activation C&W attack is less sensitive to vanishing gradient problems.
-However, we notice that, by considering only two logits at time, and neglecting the effect of the perturbation on the other logits, this approach is particularly suited to the one-hot case and is highly suboptimum when other output encoding schemes are used.
-3.2 Attacks against DNNs with output encoding
-The adversarial attack algorithms described so far focus on models based on one-hot encoding.
+However, by considering only two logits at time, and neglecting the effect of the perturbation on the other logits, this approach is particularly suited to the one-hot case and is highly suboptimum when other output encoding schemes are used.
+IV-B Attacks against DNNs with output encoding
+The adversarial attack algorithm and methods described above focus on models based on one-hot encoding.
 This is a favorable scenario for the attacker, who only needs to focus on increasing the output score (or logit) of the target class $t$.
@@
 This prevents the application of the basic adversarial algorithms described in the previous section.
-The problem of crafting adversarial examples against multi-label classifiers has been addressed in [10], where the C&W and the DeepFool attacks have been extended to work in a multi-label setting.
-The extension of C&W attack, referred to as ML-C&W, works as follows: for every output node, a hinge loss term similar to the one in Eq.
-(3.1) is considered in the minimization.
+As mentioned in Section II the problem of crafting adversarial examples against multi-label classifiers has been addressed in [9], where the C&W and the DeepFool attacks have been extended to work in a multi-label setting.
+The extension of C&W attack, hereafter referred to as ML-C&W, works as follows: for every output node, a hinge loss term similar to the one in Eq.
+(IV-A) is considered in the minimization.
 More specifically, the following term is added to the minimization: $\sum_{i=0}^{n}\max(0,\gamma-c_{ti}\cdot z_{i}(x+\delta))$, where $c_{t}$ is the target label vector and $\gamma$ is a confidence parameter.
@@
 However, by working at the logits level, the C&W attack is general and can also be extended to work with different output encoding schemes (as discussed below).
-An extension of the DeepFool attack, referred to as ML-DF, has also been proposed in [10].
-Instead of targeting an objective function minimizing the distortion, the method looks for the minimum perturbation that allows to enter the target region under a linear assumption on the model behavior.
+The extension of the DeepFool attack (ML-DF) works as follows: instead of targeting an objective function minimizing the distortion, ML-DF looks for the minimum perturbation that allows to enter the target region under a linear assumption on the model behavior.
 ML-DF works at the output score level, and hence is more prone to the vanishing gradient problem than ML-C&W.
-Given the threshold vector ${\rho}=({\rho}_{1},{\rho}_{2},..,{\rho}_{l})^{T}$ for the score level output ($\rho_{i}=0.5$ when $f_{i}\in[0,1]$, 00 when $f_{i}\in[-1,1]$), a vector with the distance to the boundary is derived and used to compute the perturbation.
+Given the threshold vector ${\rho}=({\rho}_{1},{\rho}_{2},..,{\rho}_{l})^{T}$ for the score level output ($\rho_{i}=0.5$ when $f_{i}\in[0,1]$, 0 when $f_{i}\in[-1,1]$), a vector with the distance to the boundary is derived and used to compute the perturbation.
 The perturbation obtained in this way is applied to the image.
-ML-DF applies the above procedure iteratively until the attack succeeds or a maximum number of iterations is reached.
+ML-DF applies the above procedure iteratively until the attack succeeds or the maximum number of iterations is reached.
 ML-DF is a greedy method and often leads to very large distortions; Moreover, the performance obtained are always inferior with respect to ML-C&W [9, 10].
-Inspired by ML-DF, another approach to implement an adversarial attack in the multi-label case, named Multi-Label Attack via Linear programming (MLA-LP), has been proposed in [11].
+MLA-LP, which also implements an adversarial attack for the multi-label case, works as follows.
 By assuming that the loss changes are linear for small distortions, a simplified formulation of the attack is solved to minimize the $L_{\infty}$ distortion introduced in the image subject to a constraint on the loss value, requiring that the final loss is lower than the threshold value.
-Thanks to the adoption of the $L_{\infty}$ distortion, the problem can be easily transformed into a linear programming problem, for which several solvers exist.
+Thanks to the adoption of the $L_{\infty}$ distortion, the problem can be easily solved by linear programming methods.
 Similarly to ML-DF, MLA-LP works at the final score output level.
-By considering the loss function instead of the model output function, MLA-LP tends to modify less the confidences of non-attacked labels with respect to ML-DF, with a lower distortion for the resulting attacked image.
+By considering the loss function instead of the model output function, MLA-LP tends to modify less the confidences of non-attacked labels with respect to ML-DF, with a lower distortion of the attacked image.
 When the output coding scheme is based on ECOC, an adversarial attack can be carried out by extending to this setting the C&W method.
@@
 Formally, the optimization problem solved in [20] is defined as:
-$\min_{\delta}$ $\hskip 5.69046pt\left(||\delta||_{2}^{2}-\lambda\cdot\min_{i}(c_{ti}\cdot z_{i}(x+\delta),\eta)\right),$ (6)
+[…]
 where $\eta$ is a constant parameter setting a confidence threshold for the attack.
@@
 A straightforward, highly suboptimum, way to apply these methods to the ECOC case is by setting $\rho=(0,0,...,0)^{T}$, as if all the codewords were possible, and performing the decoding element-wise.
-3.2.1 The LOTS attack
+IV-B1 The LOTS attack
 A completely different kind of attack, which can be applied to any network regardless of the output encoding scheme is the Layerwise Origin-Target Synthesis (LOTS) attack [21].
 LOTS generates the adversarial example by modifying the input sample so that its representation in the feature space is as similar as possible to that of a given target sample.
-Formally, given the internal representation $f^{-j}(x)$ of an image $x$ at layer $L-j$ and a target internal representation $f^{-j}(x_{t})$, LOTS attempts to minimize the term:
-$||f^{-j}(x^{t})-f^{-j}(x)||^{2}.$ (7)
+Formally, given the internal representation $f^{-j}(x)$ of an image $x$ at layer LjL-j and a target internal representation $f^{-j}(x_{t})$, LOTS attempts to minimize the term:
+$f^{-j}(x^{t})-f^{-j}(x)||^{2}.$ (6)
 At each iteration, the algorithm updates the input sample as
 $x^{(i)}=x^{(i-1)}-\frac{\nu^{-j}(x,x^{t})}{\max_{x}(|\nu^{-j}(x,x^{t})|)}$ (7)
-where $\nu^{-j}(x,x^{t})=\nabla_{x}\left(||f^{-j}(x^{t})-f^{-j}(x)||^{2}\right)$, and ||$\cdot$|| ⋅ | is applied element-wise, until the Euclidean distance between $f^{-l}(x)$ and the target is smaller than a predefined threshold.
-The final perturbed image $x^{adv}$ obtained in this way has an internal representation at layer $L-j$ that mimics that of the target sample $x^{t}$.
-In the LOTS algorithm, the target point $x^{t}$ is any point belonging to the target class of the attack.
-We observe that in general such $x^{t}$ may not be available, e.g.
-in the multilabel case when the attacker targets an arbitrary labels’ combination.
+[…]
 When applied to the logits level ($f^{-1}$), LOTS simultaneously modifies all the logits.
 However, by focusing on the minimization of the Euclidean loss, the distortion introduced in the input space is not controlled.
-3.2.2 JMA and prior art
-Below we discuss the relationship and advantages of JMA with respect to the state-of-the-art attacks described above.
+IV-B2 JMA and prior art
 In contrast to LOTS [21], JMA defines the target point in the feature space that minimizes perturbation necessary to move the feature representation of the input sample to the target point.
 It does so, by relying on the Mahalanobis distance induced by the Jacobian matrix of the neural network function so to take into account the different effort, in terms of input perturbation, required to move the sample in different directions.
-The ML-DF and MLA-LP methods [10, 11] are also based on the Jacobian matrix.
-However, in [10], the Jacobian matrix is used to implement a greedy algorithm, while in [11] is used to solve a suboptimum formulation of the attack.
-In our case, instead, we are able to solve the original formulation of the attack by Szegedy et al.
-under a first-order approximation.
+ML-DF and MLA-LP [9, 10] are also based on the Jacobian matrix, however, in [9], the Jacobian matrix is used to implement a greedy algorithm, while in [10] is used to solve a suboptimum formulation of the attack.
+With JMA, instead, the original formulation of the attack by Szegedy et al.
+is solved under a first-order approximation.
 With such an approximation, in fact, the solution of the attack problem can be reduced to the solution of a constrained quadratic programming problem (where the objective function is a Mahalanobis distance term induced by the Jacobian).
-Moreover, our algorithm is a general one and can work either at the output level or at the logits level, regardless of the output encoding scheme used by the network.
-4 The JMA attack
-As mentioned in the introduction, to avoid vanishing gradient problems, we look for a method that can be applied at the logits level.
-As we said, while modifying one or two logits at a time allows to carry out a close-to-optimal adversarial attack in the one-hot encoding case [7], in the more general case of DNNs based on output encoding, this approach is highly suboptimum.
+Moreover, our algorithm is a general one and can work either at the output level or at the logits level, regardless of the output encoding scheme used.
+V The JMA attack
+As mentioned in Section I, our method is designed to operate at the logits level.
+In fact, while modifying one or two logits at a time allows to carry out a close-to-optimal adversarial attack in the one-hot encoding case [6], in the more general case of DNNs based on output encoding, this approach is highly suboptimum.
 In this scenario, in addition to considering all the logits simultaneously, as done in the ML-C&W attack, the correlation among the logits and the effort required in the input space to move the input sample in a given direction (in the logit or feature space) must be considered.
-The JMA attack tackles the above problems and finds the optimal adversarial image, by solving the optimization problem in Eq.
+JMA tackles the above problems and finds the optimal adversarial image, by solving the optimization problem in Eq.
 (3), under a first-order approximation of the behavior of the network function.
-In the following, we look at the output of the neural network function $f$  [^3] For ease of notation, in the following, we consider directly the model output $f(x)$, however the same the arguments cam be applied at the logit level.
+In the following, we look at the output of the neural network function $f$  [^2] For ease of notation, in the following, the model output is considered, however the same procedure can be applied at the logit level.
 as a generic point in $\mathbb{R}^{n}$.
@@
 Let $x_{0}$ be the to-be-attacked image and let us denote with $\delta$ the perturbation applied by the attacker.
-Let $r=f(x_{0})+d$.
+Let $r=f(x_{0})+d$ denote a generic point in the output space.
 The goal of the attacker is to determine the displacement $d$, in such a way that $r$ lies inside the decision region of the target class, and for which the Euclidean norm of the perturbation $\delta(d)$ necessary to reach $r$ is minimum, that is:
-$\min_{d:\phi(f(x_{0})+d)=t}\left[\min_{\begin{subarray}{c}\delta:f(x_{0}+\delta)-f(x_{0})=d\\ x_{0}+\delta\in[0,1]^{m}\end{subarray}}\hskip 5.69046pt||\delta(d)||_{2}\right].$ (11)
+[…]
+The proposed attack works in two steps:
+  • 1. the optimum perturbation $\delta^{*}(d)$ moving the input sample to the (generic) target point $r$ is determined;
+  • 2.
+we find the optimum target point $r^{*}$, i.e., the optimum displacement $d^{*}$ (where $r^{*}=f(x_{0})+d^{*}$) by solving a constrained quadratic programming problem.
+The resulting $\delta^{*}(d^{*})$ gives the minimum perturbation necessary to reach the target region.
+In the following, we first rewrite the minimization in a more convenient way, then we solve Step 1 and 2 under a first order approximation.
 Without loss of generality, we will assume that all the label codewords have the same norm, and that $\phi()$ applies a minimum distance decoding rule.
@@
 where $g_{ti}=(f(x_{0})^{T}c_{t}-f(x_{0})^{T}c_{i})$, and hence the optimization problem in Eq. (8) can be rewritten as:
-$\min_{\begin{subarray}{c}d:d^{T}(c_{i}-c_{t})\leq g_{ti}\\ \forall i,i=1,\cdots,l\end{subarray}}\bigg{[}\min_{\delta:f(x_{0}+\delta)-f(x_{0})=d}\hskip 5.69046pt||\delta(d)||_{2}\bigg{]}.$ (16)
+[…]
 Note that we neglected the constraint $x_{0}+\delta\in[0,1]^{m}$, trusting that if $\delta$ is small the constraint is always satisfied.
-We will then reconsider the effect of this constraint at the end of our derivation in Section 4.1.
+We will then reconsider the effect of this constraint at the end of our derivation in Section V-A.
 Under the assumption that the input perturbation is small (which should always be the case with adversarial examples), we can consider a first order approximation of the effect of the perturbation $\delta$ on the network output:
 $f(x_{0}+\delta)\simeq f(x_{0})+J_{x_{0}}\delta,$ (12)
+where $J_{x_{0}}$ denotes the Jacobian matrix of $f$ in $x_{0}$333ff can describe the output of the network or the logits (strictly speaking, in the latter case, $f$ should be replaced by $f^{-1}$):
+$J_{x}=\nabla_{x}f(x)|_{x_{0}}=\left[\frac{\partial f_{i}(x_{0})}{\partial x_{j}}\right]_{i=1,\cdots,n,j=1,\cdots m}.$ (13)
+We now show that the minimization in Eq. (11) can be solved under the first order approximation in Eq. (12). In particular, Step 1 corresponds to solving the following problem:
 […]
-$J_{x}=\nabla_{x}f(x)|_{x_{0}}=\left[\frac{\partial f_{i}(x_{0})}{\partial x_{j}}\right]_{i=1,\cdots,n,j=1,\cdots m}.$ (18)
-We now show that, under the first order approximation in Eq.
-(17), we can solve the minimization in Eq.
-(16), by applying the following two steps: i) we first determine the optimum perturbation moving the input sample to a given target point $r$; ii) then, we find the optimum target point $r^{*}$ by solving a constrained quadratic programming problem.
-Addressing i) corresponds to solving the following problem:
-$\delta^{*}(d)=\arg\min_{\delta:f(x+\delta)-f(x_{0})=d}||\delta||_{2},$ (19)
 that, under the linear approximation, becomes [^4] To keep the notation light, here and in the sequel, we omit the subscript of the Jacobian matrix.
@@
 Proposition 1.
-Given that $n<m$, under the assumption that $J$ has full rank, the solution of the minimization in Eq. (20) is:
+Since $n<m$, under the assumption that $J$ has full rank [^5] We verified this assumption experimentally as discussed in Section VI-D., the solution of the minimization in Eq. (15) is:
 ${\delta}^{*}(d)=J^{T}(JJ^{T})^{-1}d.$ (16)
@@
 From Proposition 1, we can express the norm of the optimum perturbation as follows
-$||\delta^{*}(d)||^{2}=$ $\delta^{*T}\delta^{*}=(J^{T}{(JJ^{T})}^{-1}d)^{T}(J^{T}{(JJ^{T})}^{-1}d)$
+[…]
 $=$ $d^{T}(J^{T}{(JJ^{T})}^{-1})^{T}(J^{T}{(JJ^{T})}^{-1}d)$
@@
 where in the last equality we have exploited the symmetry of $(JJ^{T})^{-1}$ (being the inverse of a symmetric matrix $(JJ^{T})$), and then ${((JJ^{T})^{-1})}^{T}=(JJ^{T})^{-1}$.
-Given the above, the optimal target point $r^{*}$, and hence the optimal distance term $d^{*}$, is determined by solving the following minimization:
+The optimal target point $r^{*}$, and hence the optimal distance term $d^{*}$, is then determined (Step 2) by solving the following minimization:
 $\min_{d:d^{T}(c_{i}-c_{t})\leq g_{ti},\forall i\neq t}$ $\hskip 5.69046ptd^{T}{{(JJ^{T})}^{-1}}d.$ (20)
@@
 Figure 1: Illustration of the intuition behind the formulation in (20).
-While $r^{\prime}$ is the point in the target region closest to $f(x_{0})$ in the Euclidean norm, $r^{*}$ is the point that can be reached with the minimum distortion in the input space.
+The point in the target region closest to $f(x_{0})$ in the Euclidean norm is $r^{\prime}$.
+Then, $r^{\prime}=f(x_{0})+d^{\prime}$, where $d^{\prime}$ is the minimum norm displacement necessary to reach the target region.
+Instead, the point that can be reached with the minimum distortion in the input space, i.e., the point which minimizes $\delta$, is $r^{*}$ (this point corresponds to a displacement $d^{*}$ having larger norm than $d^{\prime}$).
+$r^{*}$ is the optimal target point that the attack wants to reach.
 To solve (20), we find convenient to rewrite it as follows:
 $\min_{d:Ad\leq b}\frac{1}{2}d^{T}{{(JJ^{T})}^{-1}}d,$ (21)
-where $A=[(c_{1}-c_{t})^{T}\dots(c_{l}-c_{t})^{T}]$ has size $(l-1)\times n$, and $b$ is an $l$-long vector defined as $b_{i}=g_{ti}$. We observe that when $l>n+1$ the system is overdetermined.
+where $A=[(c_{1}-c_{t})^{T}\dots(c_{l}-c_{t})^{T}]$ has size $(l-1)\times n$, and $b$ is an (l1)(l-1)-long vector defined as $b_{i}=g_{ti}$. We observe that when $l>n+1$ the system is overdetermined.
 Theorem 1.
@@
 $\lambda\geq 0,$ (25)
-where $\lambda$ is a column vector with $k-1$ entries.
-By rewriting the objective function as $-\frac{1}{2}d^{T}{{(JJ^{T})}^{-1}}d+d^{T}({{(JJ^{T})}^{-1}}d+A^{T}\lambda)-b^{T}\lambda$ and after some algebra, the optimization problem can be rephrased as:
+where $\lambda$ is a column vector with l1l-1 entries.
+By rewriting the objective function as $\frac{1}{2}d^{T}{{(JJ^{T})}^{-1}}d+d^{T}({{(JJ^{T})}^{-1}}d+A^{T}\lambda)-b^{T}\lambda$ and after some algebra, the optimization problem can be rephrased as:
 $\min_{d,\lambda}\hskip 5.69046pt$ $\frac{1}{2}d^{T}{{(JJ^{T})}^{-1}}d+b^{T}\lambda$
@@
 In particular, an easy-to-implement algorithm, whose complexity grows linearly with the number of label vectors, named Sequential Coordinate-Wise algorithm, has been proposed in [30].
+Theorem 1 identifies the minimum perturbation necessary to enter the target region $\delta^{*}(d^{*})$ (simply indicated as $\delta^{*}$ in the following).
+Then the adversarial image is computed as $x_{adv}=x+\delta^{*}$.
 Note that in the multi-label case, the number of label vectors grows exponentially with $n$ (since $l=2^{n}$).
-However, in this case, the problem can be significantly simplified, as discussed in Section 4.2.
-4.1 Iterative formulation of JMA
+However, in this case, the problem can be significantly simplified, as discussed in Section V-B.
+V-A Iterative formulation of JMA
 In principle, JMA should produce an adversarial example in one-shot.
-In practice, however, this is not always the case, due to the fact that the linear assumption of $f$ holds only in a small neighborhood of the input $x$, hence the assumption may not be met when the distortion necessary to attack the image becomes larger.
+In practice, however, this is not always the case, due to the fact that the linear assumption of $f$ holds only in a small neighborhood of the input $x$, hence the assumption may not be met when the distortion necessary to attack the image is larger.
 In order to mitigate this problem, we consider the following iterative version of JMA:
   • When $\delta^{*}$ does not bring $x$ into the target region, that is, $\phi(x+\delta^{*})\neq t$, we update the input in the direction given by the perturbation considering a small step $\epsilon$, thus remaining in the close vicinity of $x$.
-  • When the target class is reached, a binary search is performed between the perturbed sample $x+\delta^{*}$ and the original input $x$, and the adversarial example resulting in the smallest perturbation is taken as the final output of the attack.
-Finally, in practice, after that the optimal perturbation $\delta^{*}$ is superimposed to $x$, a clipping operation is performed to be sure that the solution remains in the $[0,1]^{m}$ range.
-Regarding the full rank of the Jacobian, this assumption has been verified experimentally and discussed in Section 5.4.
-The iterative JMA attack resulting from the application of the above steps is described in Algorithm 1. In the algorithm, the maximum number of iterations (updates) is set to $n_{it,max}$.
-Algorithm 1 Iterative Jacobian-induced MAhalanobis distance (JMA) attack
+  • When the target class is reached, a binary search is performed between the perturbed sample $x+\delta^{*}$ and the original input $x$, and the adversarial example resulting in the smallest perturbation is considered.
+Finally, in practice, after that the optimal perturbation $\delta^{*}$ is superimposed to $x$, a clipping operation is performed to be sure that the solution remains in the $0,1]^{m}$ range.
+The iterative version of JMA resulting from the application of the above steps is described in Algorithm 1. In the algorithm, the maximum number of iterations (updates) is set to $n_{it,max}$.
+Algorithm 1 Jacobian-induced Mahalanobis-distance attack
 0:   max no. of iterations $n_{it,max}$, to be attacked image $x$, target class $t$ (target label vector $c_{t}$), updating step size $\epsilon$
@@
 3:  calculate the Jacobian matrix by backward propagation: $J=\nabla_{x}f(x)$
-4:     calculate matrix $A$ and vector $b$
+4:  calculate matrix AA and vector $b$
 5:  calculate $\lambda^{*}$ in Eq.(23) via the sequential Coordinate-Wise algorithm in [30]
@@
 15:  else
-16:        $x=(x-\delta^{*})+\epsilon\frac{\delta^{*}}{||\delta^{*}||}$
+[…]
 17:  end if
 18:end for
-4.2 Simplified formulation for the multi-label case
+V-B Simplified formulation for the multi-label case
 In most cases, the number of classes $l$, determining the number of rows of matrix $A$ in Eq.
@@
 Therefore, the NNLS problem can be solved by means of the sequential Coordinate-Wise algorithm in [30] with complexity $O(n)$ instead of $O(2^{n})$ as in the original formulation.
-5 Experimental Setting
+VI Experimental Setting
 To evaluate the performance of JMA, we have run several experiments addressing different classification scenarios, including ECOC-based classification [8] and multi-label classification.
-Although less significant for our study, we have also run some experiments in the one-hot encoding scenario.
-In all the cases, the JMA attack is applied at the logits level.
-We implemented the JMA attack by using Python 3.6.9 via the Keras 2.3.1 API.
+Although less significant for this study, we have also run some experiments in the one-hot encoding scenario.
+In all the cases, JMA is applied at the logits level.
+We implemented JMA by using Python 3.6.9 via the Keras 2.3.1 API.
 We run the experiments by using an NVIDIA GeForce RTX 2080 Ti GPU.
-The code, as well as the trained models and all the information for the reproducibility of the experiments are publicly available at https://github.com/guowei-cn/JMA–A-General-Close-to-Optimal-Targeted-Adversarial-Attack-with-Improved-Efficiency.git.
-In the following sections, we describe the classification tasks considered in the experiments, present the evaluation methodology and comparison analysis, and detail the experimental setting.
-5.1 Classification networks and settings
-5.1.1 One-hot encoding
-To assess the performance of JMA attack in the one-hot encoding scenario, we considered traffic sign classification on the German Traffic Sign Recognition Benchmark (GTSRB) dataset [25], with a VGG16 architecture.
-The error rate of the trained models on clean images is 0.005.
-5.1.2 ECOC-based classifier
-We considered the ECOC framework for three image classification tasks, namely MNIST classification [26], CIFAR-10 classification [27], and GTSRB classification [25].
+The code, as well as the trained models and the information for the reproducibility of the experiments are publicly available at https://github.com/guowei-cn/JMA–A-General-Close-to-Optimal-Targeted-Adversarial-Attack-with-Improved-Efficiency.git.
+In the following sections, we describe the classification tasks considered in the experiments, present the evaluation and comparison methodologies, and detail the experimental setting.
+VI-A Classification networks and settings
+VI-A1 One-hot encoding
+To assess the performance of JMA in the single-label classification scenario, we considered the task of traffic sign classification on the German Traffic Sign Recognition Benchmark (GTSRB) dataset [31], with a VGG16 architecture.
+In this scenario, the accuracy of the the trained model on clean images is 0.995.
+Moreover, to validate the effectiveness of the JMA algorithm in a challenging classification scenario with a large number of classes, we defined a new classification task by selecting 2,000 classes from the most populated categories in the ImageNet21K dataset [32] (referred to as ImageNet2K in the following).
+A ViT-B/16 network pretrained on ImageNet21K and fine-tuned on a training subset of ImageNet2K was used for this task.
+The accuracy on clean images is 0.66.
+VI-A2 ECOC-based classifier
+We considered the ECOC framework for three image classification tasks, namely MNIST [33], CIFAR-10 [34], and GTSRB [31].
 We implemented the ECOC scheme considering an ensemble of networks, each one outputting one bit [^6] According to [8], resorting to an ensemble of networks permits to achieve better robustness against attacks..
@@
 For the GTSRB case, we set $l=32$, by selecting the classes with more examples, and used an Hadamart code with $n=32$.
-Following the original ECOC design, we considered the VGG16 architecture [28] as baseline, with the first 6 convolutional layers shared by all the networks of the ensemble (shared bottom), and the remaining 10 layers (the last 8 convolutional layers and the 2 fully connected layers) that have been trained separately for each ensemble branch.
-The weights of the Image-Net pre-trained model are used for the shared bottom.
-5.1.3 Multi-label classification
-For the multi-label experiments, we considered the model architecture adopted in [10] that is, a standard InceptionV3 [29].
+Following the original ECOC design, we considered the VGG16 architecture [35] as baseline, with the first 6 convolutional layers shared by all the networks of the ensemble (shared bottom), and the remaining 10 layers (the last 8 convolutional layers and the 2 fully connected layers) trained separately for each ensemble branch.
+The weights of the ImageNet pre-trained model are used for the shared bottom.
+VI-A3 Multi-label classification
+We primarily tested the performance of JMA in a multi-label setting by using the VOC2012 dataset [36], which is a benchmark dataset adopted in several multi-class classification works [24, 23].
+This dataset has been used to train models to recognize object inside scenes.
+Specifically, the 11,530 images of the dataset contain objects from 20 classes.
+Then, $n=20$ and the number of possible label vectors is $2^{n}\approx 10^{6}$.
+The dataset is split into training, validation, and testing subsets, with proportion 6:2:2.
+We considered the same model architecture adopted in [9] that is, a standard InceptionV3 [37].
 The categorical hinge loss was used to train the model, with a standard Adam optimizer with $lr=10^{-4}$ and batch size 64.
-TABLE I: Types of objects in the VOC2012 dataset.
-Category Object name
-Person person
-Animal bird, cat, cow, dog, horse, sheep
-Vehicle aeroplane, bicycle, boat, bus, car, motorbike, train
-Indoor bottle, chair, dining table, potted plant, sofa, tv/monitor
-The model was trained on the VOC2012 dataset [30], which is a benchmark dataset adopted in several multi-class classification works [14, 31].
-This dataset has been used to train models to recognize object inside scenes.
-Specifically, the 11,530 images of the dataset contain objects from 20 possible object classes, see Table I (the objects are classified into four categories).
-Then, $n=20$ and the number of possible label vectors is $2^{n}\approx 10^{6}$.
-All datasets are split into training, validation, and testing subsets, with proportion 6:2:2.
 The performance of the network in this case are measured in terms of mean average precision (mAP), that is, the mean of the average precision (AP) for all the classes, where AP is a measure of the area under the precision-recall curve.
 According to our experiments, the mAP of our model is 0.93, which is aligned with [9].
-5.2 Comparison with the state-of-the-art
-We have compared the performance of JMA with those of the most relevant state-of-the-art attacks for each classification scenario.
-In all the cases, we used the codes made available by the authors in their repositories.
-For the multi-label case, we considered ML-C&W [10], MLA-LP [11] and LOTS [18] (we have not considered ML-DF as its performance are always inferior to ML-C&W, see [10, 11]).
-As described in Section 3, LOTS can be applied to any internal layer of the network.
-In our experiments, for a fair comparison, we have applied LOTS at the logits layer.
-By following [18], for every target label vector, we randomly selected 20 images with that label vector, and averaged the logits to get the target output vector (for target label vectors for which the total number of images available is lower than 20, we used all the images).
+To better show the generality of the proposed attack, we also trained additional models using different and more modern architectures, namely, ResNet50 [38], ViT-B/16 [39], and CLIP+MLP [40] , and ran attacks against them The results regarding these networks are reported in Section VII-D.
+We further validated the effectiveness of JMA on multi-label classification by running some additional experiments on more complex datasets, namely MS-COCO [41] and NUS-WIDE [42] (see again Section VII-D).
+VI-B Comparison with the state-of-the-art
+We have compared the performance of JMA with the most relevant state-of-the-art attacks for the various classification scenarios.
+In all the cases, we used the code made available by the authors in public repositories.
+For the multi-label case, we considered ML-C&W [9], MLA-LP [10] , LOTS [21] and the method in [11], referred to as SemA-ML in the following.
+We did not consider ML-DF, since its performance are always inferior to those of ML-C&W [9, 10])
+Regarding LOTS, as described in Section IV, it can be applied to any internal layer of the network.
+In the experiments, for a fair comparison, we have applied LOTS at the logits layer.
+By following [21], for every target label vector, we randomly selected 20 images with that label vector, and averaged the logits to get the target output vector (for target label vectors for which the total number of images available is lower than 20, we used all of them).
 We stress that, in the way it works, LOTS can not target an arbitrary labels’ combination, since examples for that combination may not be available.
-For the case of ECOC classification, the comparison is carried out against C&W, which is also the attack considered in [9], and [17] (ECOC Attack) specialized to work with the ECOC model [17].
-Although suboptimal in this case, we also considered ML-C&W and MLA-LP, by applying the latter as discussed in Section 3.2.
-However, given the complexity of the attack in the ECOC scenario and the high sub-optimality of MLA-LP, the performance of this algorithm are extremely poor, and no image can be attacked by using this method.
-Finally, the performance of LOTS attack are also assessed.
-In this case, the target output vectors are computed by averaging on 50 images (when the number of available images is lower than 50, we averaged all of them).
+For the case of ECOC-based classification, the comparison is carried out against C&W, which is also the attack considered in [8], and [20] (ECOC Attack) specialized to work with the ECOC model [20].
+Although suboptimal in this case, we also considered ML-C&W and MLA-LP, by applying the latter as discussed in Section IV-B.
+However, given the complexity of the attack in the ECOC scenario and the high sub-optimality of MLA-LP, the performance of this attack are extremely poor, and no image can be attacked by using it.
+Finally, the performance of LOTS are also assessed.
+In this case, the target output vectors are computed by averaging 50 images (when the number of available images is lower than 50, we averaged all of them).
 For the one-hot encoding scenario, we considered the original C&W algorithm [6], which is one of the best performing white-box attack working at the logits level.
-5.3 Evaluation methodology
-To test the attacks in the various scenarios, we proceeded as follows: we randomly picked 200 images from the test set (among those that are correctly classified by the network model); then, for each image, we randomly picked a target label vector, different from the true label vector.
+We did not consider attacks like I-FGSM, PGD, and also AutoPGD and AutoAttack, as they cannot be applied to frameworks different than single-label classification (see discussion in Section II), e.g.
+to ECOC and in particular to the multi-label case, for which they would need a suitable extension via the definition of a proper loss function.
+VI-C Evaluation methodology
+To test the attacks in the various scenarios, we proceeded as follows: we randomly picked 200 images from the test set, among those that are correctly classified by the network model (in the complex multi-label classification tasks of MS-COCO and NUS-WIDE, some label errors always occur, hence the attacked samples are chosen at random); then, for each image, we randomly picked a target label vector, different from the true label vector.
 For multi-label classification, in principle, there can be - and in fact there are - combinations of labels that are not represented in the training set.
-it is also possible that some of these combinations do not correspond to valid label vectors, as some labels may not be simultaneously 1 (or 0).
-Since LOTS requires the availability of samples belonging to the target label vector, for ease of comparison, in our experiments, we considered label vectors that appear in the training set as target label vectors.
+it is also possible that some of these combinations do not correspond to valid label vectors.
+Since LOTS requires the availability of samples belonging to the target label vector, for ease of comparison, in the experiments, we considered label vectors that appear in the training set as target label vectors.
 For the other methods, some tests have also been carried out in the more challenging case where the target label vectors correspond to randomly chosen target vectors.
@@
 We measured the performance of the attacks in terms of Attack Success Rate (ASR), namely, the percentage of generated adversarial examples that are assigned to the target class.
-To measure the quality of the adversarial image, we considered the Peak Signal-Noise Ratio (PSNR), characterizing the distortion introduced by the attack.
-The PSNR is evaluated as $20\log_{10}(255\cdot\sqrt{H\times W}/||\delta||_{2})$, where $H\times W$ is the size of the image [^7] Given that the attacks are multi-step and iterative, and different iterations are necessary for the various algorithms, we can not use $\epsilon$ as an indication of the distortion..
-For the tests with random choice of the target vector (for a prescribed number of flipped labels), we also report the average percentage of labels/bits successfully modified by the attack, indicated as bASR.
+To measure the quality of the adversarial image, we considered the Mean Square Error (MSE) of the attack, evaluated as MSE $\delta||_{2}/\sqrt{H\times W\times 3}$, where $H\times W\times 3$ is the size of the image (image values are in the [0,1] range).
+For the tests with random choice of the target vector, we also report the average percentage of labels/bits successfully modified by the attack, indicated as bASR.
 The computational complexity is measured by providing the average number of iterations required by the attack, and the time - in seconds - necessary to run it.
-In the following, we let $n_{it}$ denote the number of iterations necessary to run an attack, or the number of attack updates.
+In the following, we let $n_{it}$ denote the number of iterations necessary to run an attack, that is, the number of attack updates.
 The average number of iterations is denoted by $\bar{n}_{it}$.
-5.4 Setting the Attack Parameters
+VI-D Attack Parameters Setting
 The tunable parameters of JMA are the maximum number of iterations (${n}_{it,max}$) and the step size $\epsilon$.
-The iteration parameter takes also into account the number of ’for’ loops and the number of steps of the binary search in the final iteration (see Algorithm 1), the total number of updates of JMA, then, is ${n}_{it}=(v-1)+n_{bs}$, where $v$ is the number of ’for’ loops of the algorithm ($v<{n}_{it,max}$, see Algorithm 1) and $n_{bs}$ is the number of steps of the binary search.
+The maximum number of iterations takes also into account the number of ‘for’ loops and the number of steps of the binary search in the final iteration (see Algorithm 1), the total number of updates of JMA, then, is ${n}_{it}=(v-1)+n_{bs}$, where $v$ is the number of ‘for’ loops of the algorithm ($v<{n}_{it,max}$, see Algorithm 1) and $n_{bs}$ is the number of steps of the binary search.
 $n_{bs}$ is set to 6 in our the experiments.
 For C&W, ML-C&W and ECOC attack, the parameters are the initial $\lambda$, the number of steps of the binary search $n_{bs}$, and the maximum number of iterations $n_{it,max}$.
-All these attacks carry out a binary search over the parameters $\lambda$ at every iteration to find the best value.
-Then, $n_{it,max}$ provides a constraint on $n_{bs}\cdot b$, where $b$ is the number of iterations of the optimization algorithm.
-As for their setting, we consider $n_{bs}=5$ and $10$, and various values of $\lambda$.
-Finally, for MLA-LP and LOTS, the only parameter is the maximum number of iterations $n_{it,max}$ allowed in the gradient descent.
-In our experiments, we considered various values of $n_{it,max}$.
+For all these attacks, a binary search is carried out over the loss tradeoff parameter $\lambda$ (see Section IV-A).
+At each binary search step, iterations are run (until $n_{it,max}$) to obtain the adversarial example.
+The number of iterations that we report considers only the number of attack updates corresponding to the final $\lambda$ value used by the attack.
+Hence, in these cases, the actual complexity of the attack is better reflected by the attack time.
+As for the setting of these parameters, we consider $n_{bs}=5$ and $10$, and various values of the initial $\lambda$.
+For SemA-ML, we set $n_{it,max}=300$ following the official code, with the other settings left as the default.
+For MLA-LP and LOTS, the only parameter is the maximum number of iterations $n_{it,max}$ allowed in the gradient descent.
+In the experiments, we considered various values of $n_{it,max}$.
 For MLA-LP, we set $n_{it,max}=100$ for the multi-label classification case, and raised it up to $3000$ in the ECOC case when the attack is more difficult.
 For LOTS, we set $n_{it,max}=2000$, namely the maximum number considered in [21], maximizing the chances to find an adversarial example.
-Given that the time complexity of LOTS is limited compared to the other comparison methods, in fact, the time complexity of the attack remains small also for large values of $n_{it,max}$.
+Given that the time complexity of LOTS is limited compared to the other methods, in fact, the time complexity of the attack remains small also for large values of $n_{it,max}$.
 We have verified experimentally that LOTS converges within 1000 iterations 92% of the times [^7] Convergence is determined by checking whether the new loss value is close enough to the average loss value of the last 10 iterations., validating the soundness of our choice.
-For ML-C&W, we set $n_{it,max}=1000$ as default value, and in most of the cases results are reported for the setting $(\lambda,n_{bs},n_{it,max})=(0.01,10,1000)$, that resulted in the highest ASR.
-Finally, for JMA, we generally got no benefit by increasing the number of iterations beyond 200, since most of the time JMA converges in few iterations, and 200 is more than enough to get and adversarial image in all the cases.
-For sake of simplicity and comparison, the confidence parameter is set to 0 for all the attack methods.
+For ML-C&W, we set $n_{it,max}=1000$ as default value, and in most of the cases results are reported for $(\lambda,n_{bs},n_{it,max})=(0.01,10,1000)$, that resulted in the highest ASR.
+Finally, for JMA, we generally got no benefit by increasing the number of iterations beyond 200, since most of the time JMA converges in by far less iterations, and 200 is more than enough to get and adversarial image in all the cases.
+To simplify the comparison, the confidence parameter is set to 0 for all the attacks.
 To avoid that the an attack gets stuck with some images (e.g.
@@
 In this case, an adversarial perturbation can not be found, resulting in a failure of the attack.
-6 Results
-The results we obtained in the various classification settings are reported and discussed in this section.
+VII Results
+The results we obtained in the various settings are reported in this section.
 In all the tables, $\bar{n}_{it}$ refers to the average number of iterations/updating of the attack, averaged on the successful attacks.
-The attack parameters we report are $(\lambda,n_{bs},n_{it,max})$ for C&W, ML-C&W and ECOC attack, and $(\epsilon,n_{it,max})$ for JMA.
-6.1 One-hot encoding
-As we said, the main advantage of JMA with respect to state of the art methods is obtained for classifiers that do not adopt one-hot encoding.
-In such a scenario, in fact, focusing only on the target logit or final score, and ignoring the others, as done by the most common adversarial attacks, is approximately optimal.
-Nevertheless, we verified that the JMA method is also effective in the one-hot encoding scenario.
-Table II reports the results we got on the GTSRB dataset.
-We see that JMA achieves a much higher ASR compared to C&W, getting close to 1, yet with a lower PSNR.
-This is likely due to the fact that JMA tends to modify all the output scores, while C&W directly focuses on raising the target output, which is an effective strategy when the one-hot scheme is adopted for the encoding.
-In addition, JMA can obtain an adversarial image in a much smaller number of iterations, requiring only very few seconds (2.45 sec in the best setting) to attack an image thus resulting much faster than C&W attack.
-TABLE II: Results of the attacks against an one-hot encoding classifier for the GTSRB dataset.
-Parameters ASR PSNR $\bar{n}_{it}$ Time(s)
-JMA ( 0.01, 200) 0.45 43.16 66.49 9.82
-(0.05, 200) 0.81 37.88 50.65 7.44
-(0.1, 200) 0.88 36.86 38.00 5.66
-(0.2, 200) 0.94 35.76 29.63 4.41
-(0.6,200) 0.98 33.67 16.20 2.46
-C&W (1e-4,5, 100) 0.19 46.65 85.95 45.06
-(1e-4,5,200) 0.33 45.24 171.88 90.11
-(1e-2,5,500 ) 0.55 42.36 443.18 233.45
-(1e-1,10,2000) 0.75 41.14 1661.11 875.35
-(1e-1,10,5000) 0.82 41.02 3577.5 1884.59
-6.2 ECOC-based classifiers
-The results we got for ECOC-based classification on GTSRB, CIFAR-10, and MNIST are reported in Tables III-V.
-We can see that, in all the cases, JMA achieves a much higher ASR for a similar PSNR, with a significantly lower computational complexity.
-In the case of GTSRB (Table III), the JMA attack achieves ASR=0.98 at PSNR=38.08dB with a total average number of iterations $\bar{n}_{it}=29.15$ and attack time of $9.41$ sec.
+The attack parameters reported are $(\lambda,n_{bs},n_{it,max})$ for C&W, ML-C&W and ECOC attack, $(\epsilon,n_{it,max})$ for JMA and $n_{it,max}$ for LOTS and MLA-LP.
+VII-A One-hot encoding
+The main advantage of JMA with respect to state of the art is obtained for classifiers that do not adopt one-hot encoding.
+In such a scenario, in fact, focusing only on the target logit or the final score, as done by the most common adversarial attacks, is approximately optimal.
+Nevertheless, we verified that JMA is also effective in the one-hot encoding scenario.
+Table I reports the results on GTSRB.
+We see that JMA achieves a larger ASR compared to C&W, yet with a larger MSE.
+This is due to the fact that JMA tends to modify all the output scores, while C&W directly focuses on raising the target output, which is an effective strategy in the one-hot encoding case.
+In addition, JMA can obtain an adversarial image with much lower iterations, requiring only very few seconds (2.45 sec in the best setting) to attack an image.
+The results of JMA against ImageNet2K are reported in Table II.
+We observe that, although JMA is not specifically designed for this scenario, it can successfully attack all the images in very few iterations, with an MSE similar to that obtained on GTSRB.
+The attack time is larger, due to the higher complexity of the NNLS problem.
+TABLE I: Results against one-hot encoding classification for GTSRB.
+Parameters ASR MSE $\bar{n}_{it}$ Time(s)
+JMA (0.05, 200) 0.81 1.6e-4 50.65 7.44
+(0.1, 200) 0.88 2.1e-4 38.00 5.66
+(0.2, 200) 0.94 2.6e-4 29.63 4.41
+(0.6,200) 0.98 4.3e-4 16.20 2.46
+C&W (1e-4,5, 100) 0.19 2.2e-5 85.95 45.06
+(1e-4,5,200) 0.33 3.0e-5 171.88 90.11
+(1e-2,5, 500) 0.55 5.8e-5 443.18 233.45
+(1e-1,10, 2000) 0.75 7.7e-5 1661.11 875.35
+(1e-1,10, 5000) 0.82 7.9e-5 3577.5 1884.59
+TABLE II: Results of JMA against one-hot encoding classification with a large number of classes (ImageNet2K).
+Parameters ASR MSE $\bar{n}_{it}$ Time(s)
+JMA (0.1, 200) 0.99 3.8e-6 26.59 445.61
+(0.2, 200) 1.00 6.2e-6 19.17 297.00
+(0.6, 200) 1.00 2.1e-5 13.59 178.78
+(1, 200) 1.00 4.2e-5 12.11 160.06
+VII-B ECOC-based classification
+The results for ECOC-based classification on GTSRB, CIFAR-10, and MNIST are reported in Tables III-V.
+In all the cases, JMA achieves a much higher ASR for a similar MSE, with a significantly lower computational complexity.
+In the tables, NA stands for ‘Not Applicable’, when no image can be attacked (ASR = 0) and then the corresponding metric can not be evaluated.
+Specifically, in the case of GTSRB (Table III), JMA achieves ASR=0.98 with MSE = 1.5e-4, a total average number of iterations $\bar{n}_{it}=29.15$ and attack time of $9.41$ sec.
 This represents a dramatic improvement with respect to the other methods.
-In the case of ECOC attack, for instance, the best result is ASR=0.93, reached for a similar distortion (PSNR=37.24dB), with $\bar{n}_{it}$ larger than 1000 ($\bar{n}_{it}=1089.03$), which is two orders of magnitude larger than JMA.
-We also see that the C&W method fails in this case, with the ASR being about 0.30 for comparable PSNR values.
-We remind that the C&W method is not designed to work in the ECOC case (the results are aligned with those reported in [9, 17]), and when applied to the ECOC classification scenario it looses part of its effectiveness (see discussion in Section 3).
-ML-C&W is more effective, however it can only achieve ASR=0.78 with a significantly lower PSNR (30.13dB) and at the price of a larger complexity.
-A very poor behavior of the attack is observed for LOTS and in particular MLA-LP, that is never able to attack the test images, notwithstanding the large value of $n_{it,max}$.
+In the case of ECOC attack, for instance, the best result is ASR=0.93, reached for a similar distortion (MSE = 1.9e-4), with $\bar{n}_{it}$ larger than 1000 ($\bar{n}_{it}=1089.03$), which is two orders of magnitude larger than JMA.
+We also see that the C&W method fails in this case, with the ASR being about 0.30 for comparable MSE values.
+Remind that the C&W method is not designed to work in the ECOC case (the results are aligned with those reported in [8, 20]), and when applied to the ECOC classification scenario it loses part of its effectiveness (see discussion in Section IV).
+ML-C&W is more effective, however it can only achieve ASR=0.78 with a significantly larger MSE (9.7e-4) and at the price of a larger complexity.
+A very poor behavior of the attack is observed for LOTS and in particular MLA-LP, that is never able to attack the test images in 1 hour, notwithstanding the large value of $n_{it,max}$.
 Since this method implements a greedy approach, it is not surprising that it fails in the complex ECOC classification scenario (more in general, the performance of MLA-LP are poor whenever the attack aims at changing several bits/labels - see the results on multi-label classification).
-For CIFAR-10 (Table IV), JMA achieves ASR = 1 with a PSNR of about 40dB, with very low complexity.
-C&W attack achieves a much lower ASR for a similar PSNR, with higher complexity.
-The ECOC attack can achieve ASR = 0.99 with PSNR=37.27dB in the setting with $n_{it,max}=2000$, in which case the computational complexity of the algorithm is high ($\bar{n}_{it}=1347.42$ for an average attack time of $557.86$ sec).
-Regarding LOTS, it can achieve ASR = 1, yet with a lower PSNR (35.79 dB).
-The results we got with LOTS confirm the lower complexity of this method with respect to ECOC attack and C&W.
-We also observe that, in this case, ML-C&W have poorer performance than C&W, with only 56% of the images that can be attacked with a lower PSNR, and with a very high computational cost.
-Finally, as before, MLA-LP is not able to attack the image in almost all the cases (ASR = 0.04).
+In the case of CIFAR-10 (Table IV), JMA achieves ASR = 1 with a pretty small distortion (MSE = 1.1e-4) and a very low complexity.
+C&W attack achieves a much lower ASR (ASR = 0.49) for a slightly lower MSE, with higher complexity.
+The ECOC attack can achieve ASR = 0.99 with MSE=1.9e-4 in the setting with $n_{it,max}=2000$, in which case the computational complexity of the algorithm is high ($\bar{n}_{it}=1347.42$ for an average attack time of $557.86$ sec).
+Regarding LOTS, it can achieve ASR = 1, yet with a larger MSE.
+The results with LOTS confirm the lower complexity of this method with respect to ECOC attack and C&W.
 The effectiveness of JMA is also verified in the MNIST case (see Table V).
-As it is always the case when attacking MNIST, the PSNR of the attacked images is lower since, unavoidably, the attacker has to modify more the image (this is particulary true in the targeted attack scenario).
-In this case, JMA reaches ASR = 1 with PSNR = 23.30 dB, while the ASR of ECOC attack (in the best possible setting) and LOTS is respectively 0.73 and 0.78 for a similar PSNR.
+In this case, JMA reaches ASR = 1 with MSE=2.3e-3, while the ASR of ECOC attack (in the best possible setting) and LOTS is respectively 0.73 and 0.78 for a similar MSE.
 The computational time is higher for LOTS and much higher for the ECOC attack.
-Both C&W and ML-C&W have poor performance and the ASR is lower than 0.52 for a PSNR similar to that obtained by JMA.
+Both C&W and ML-C&W have poor performance with an ASR lower than 0.52 for a MSE similar to that obtained by JMA.
 Fig. 2 reports some examples of images (successfully) attacked with the various methods for the three tasks.
@@
 The attacker’s goal is to cause a misclassification from ‘cat’ to ‘airplane’ (CIFAR-10), from speed limitation ‘80kph’ to ‘30kph’ (GTSRB), and from digit ‘1’ to ‘3’ (MNIST).
-The corresponding PSNR and time required for the attack are also provided for each case.
-TABLE III: Results against ECOC-based classification on the GTSRB dataset.
-Parameters ASR PSNR ${\scriptsize\bar{n}_{it}}$ Time(s)
-JMA (0.05, 200) 0.93 39.12 58.46 18.14
-(0.1, 200) 0.98 38.08 29.15 9.41
-(0.2, 200) 0.99 36.98 25.30 8.27
-C&W (1e-4,5,100) 0.08 49.38 65.69 51.31
-(1e-4,5,200) 0.11 48.02 97.73 97.67
-(1e-4,5,500) 0.12 47.42 148.33 235.76
-(1e-2,5,500) 0.21 42.12 194.19 221.63
-(1e-1,10,2000) 0.31 39.77 782.29 1453.85
-ECOC attack (1e-4,5,100) 0.35 43.42 71.22 87.87
-(1e-4,5,200) 0.40 44.18 109.90 305.28
-(1e-4,5,500) 0.43 44.45 175.67 332.88
-(1e-2,5,500) 0.61 40.64 309.03 501.80
-(1e-1,10,2000) 0.93 37.24 1089.03 2312.64
-ML-CW (0.01, 10, 1000) 0.78 30.13 111.02 2973.17
-LOTS $n_{it,max}=2000$ 0.40 39.21 395.33 97.52
-MLA-LP $n_{it,max}=1000$ 0 NA NA NA ($>$ 3600)
-TABLE IV: Results against ECOC-based classification on CIFAR-10 dataset.
-Parameters ASR PSNR $\bar{n}_{it}$ Time(s)
-JMA (0.05, 200) 1.00 39.39 14.78 3.14
-(0.1, 200) 1.00 38.79 6.88 1.92
-(0.5, 200) 1.00 38.20 6.10 1.20
-C&W (1e-4,5,100) 0.49 40.82 40.76 23.20
-(1e-4,5,500) 0.60 41.42 119.45 68.49
-(1e-4,10,200) 0.90 37.41 83.97 48.15
-(1e-4,10,500) 0.96 37.62 187.96 107.78
-(1e-1,10,2000) 0.99 37.87 593.09 340.09
-ECOC attack (1e-4,5,100) 0.65 39.11 83.08 50.36
-(1e-4,5,500) 0.92 38.56 337.53 180.37
-(1e-4,10,200) 0.93 37.78 176.85 252.89
-(1e-4,10,500) 0.95 37.33 383.74 557.86
-(1e-1,10,2000) 0.99 37.27 1347.42 2173.31
-ML-CW (0.01, 10, 1000) 0.565 29.86 49.57 1409.02
-LOTS $n_{it,max}=2000$ 1.00 35.79 38.85 8.61
-MLA-LP $n_{it,max}=1000$ 0.04 39.39 5 7.04
-TABLE V: Results against ECOC-based classification on MNIST dataset.
-Parameters ASR PSNR $\bar{n}_{it}$ Time(s)
-JMA (0.05,200) 0.95 24.08 48.61 196.55
-(0.1,200) 0.96 23.87 31.60 136.07
-(0.2,200) 1.00 23.40 16.92 62.36
-(0.5,200) 1.00 22.34 12.04 45.53
-(1,200) 1.00 20.32 8.69 33.58
-C&W (1e-3,10,100) 0.01 26.64 53.00 48.96
-(1e-3,10,200) 0.11 22.17 93.82 85.91
-(1e-3,10,500) 0.37 20.78 176.81 163.32
-(1e-3,10,1000) 0.42 20.87 328.05 327.64
-(1e-1,10,2000) 0.52 20.59 910.73 841.25
-ECOC attack (1e-3,10,100) 0.18 21.11 92.28 131.46
-(1e-3,10,200) 0.38 21.32 181.63 259.54
-(1e-3,10,500) 0.59 21.58 488.73 697.33
-(1e-3,10,1000) 0.69 21.87 967.56 1387.41
-(1e-1,10,2000) 0.73 22.24 1921.29 2754.74
-ML-CW (0.01, 10, 1000) 0.64 18.47 297.53 1679.72
-LOTS $n_{it,max}=2000$ 0.78 23.66 242.23 600.14
-MLA-LP $n_{it,max}=1000$ 0.00 NA NA NA (1079 s)
-MLA-LP $n_{it,max}=3000$ 0.00 NA NA NA ($>$3600 s)
-6.3 Multi-label classification
-Table VI. shows the results we got in the case of multi-label classification when the target label vector is chosen randomly among those contained in the training dataset.
-The average number of labels targeted by the attacks is 2.1999Most of the images in the VOC2012 dataset contains one or very few labeled objects, hence the label vectors have few 1’s and choosing the target label from the training dataset results in few bit changes.
-We see that both JMA and ML-C&W attack get ASR = 1.
-However, JMA has an advantage over ML-C&W both in terms of PSNR (around 3dB) and, above all, in terms of computational complexity, being approximately 64 times faster than ML-C&W.
-In contrast to what one would expect, LOTS achieves much worse performance, with ASR = 0.66 and a lower PSNR.
+The distortion and time required for the attack are also provided for each case.
+TABLE III: Results against ECOC- classification on GTSRB. NA means ‘Not Applicable’ (the metric can not be evaluated as no image can be attacked).
+Parameters ASR MSE $\scriptsize\bar{n}_{it}$} Time(s)
+JMA (0.05, 200) 0.93 1.2e-4 58.46 18.14
+(0.1, 200) 0.98 1.5e-4 29.15 9.41
+(0.2, 200) 0.99 2.0e-4 25.30 8.27
+C&W (1e-4, 5, 100) 0.08 1.2e-5 65.69 51.31
+(1e-4, 5, 200) 0.11 1.6e-5 97.73 97.67
+(1e-4, 5, 500) 0.12 1.8e-5 148.33 235.76
+(1e-2, 5, 500) 0.21 6.1e-5 194.19 221.63
+(1e-1, 10, 2000) 0.31 1.1e-4 782.29 1453.85
+ECOC
+Attack
+(1e-4, 5, 100) 0.35 4.5e-5 71.22 87.87
+(1e-4, 5, 200) 0.40 3.8e-5 109.90 305.28
+(1e-4, 5, 500) 0.43 3.6e-5 175.67 332.88
+(1e-2, 5, 500) 0.61 8.6e-5 309.03 501.80
+(1e-1, 10, 2000) 0.93 1.9e-4 1089.03 2312.64
+ML-CW (0.01, 10, 1000) 0.78 9.7e-4 111.02 2973.17
+LOTS $2000$ 0.40 1.2e-4 395.33 97.52
+MLA-LP $1000$ 0.00 NA NA (>> 1h)
+TABLE IV: Results against ECOC classification on CIFAR10.
+Parameters ASR MSE $\bar{n}_{it}$ Time(s)
+JMA (0.05, 200) 1.00 1.1e-4 14.78 3.14
+(0.1, 200) 1.00 1.3e-4 6.88 1.92
+(0.5, 200) 1.00 1.5e-4 6.10 1.20
+C&W (1e-4, 5, 100) 0.49 8.3e-5 40.76 23.20
+(1e-4, 5, 500) 0.60 7.2e-5 119.45 68.49
+(1e-4, 10, 500) 0.96 1.7e-4 187.96 107.78
+(1e-1, 10, 2000) 0.99 1.6e-4 593.09 340.09
+ECOC
+Attack
+(1e-4, 5, 100) 0.65 1.2e-4 83.08 50.36
+(1e-4, 5, 500) 0.92 1.4e-4 337.53 180.37
+(1e-4, 10, 500) 0.95 1.8e-4 383.74 557.86
+(1e-1, 10, 2000) 0.99 1.9e-4 1347.42 2173.31
+ML-CW (0.01, 10, 1000) 0.565 1.0e-3 49.57 1409.02
+LOTS $2000$ 1.00 2.6e-4 38.85 8.61
+MLA-LP $1000$ 0.04 1.1e-4 5 7.04
+TABLE V: Results against ECOC classification on MNIST.
+Parameters ASR MSE $\bar{n}_{it}$ Time(s)
+JMA (0.05, 200) 0.95 3.9e-3 48.61 196.55
+(0.5, 200) 1.00 5.8e-3 12.04 45.53
+(1, 200) 1.00 9.3e-3 8.69 33.58
+C&W (1e-3, 10, 100) 0.01 2.2e-3 53.00 48.96
+(1e-3, 10, 500) 0.37 8.4e-3 176.81 163.32
+(1e-3, 10, 1000) 0.42 8.2e-3 328.05 327.64
+(1e-1, 10, 2000) 0.52 8.7e-3 910.73 841.25
+ECOC
+Attack
+(1e-3,10,100) 0.18 7.7e-3 92.28 131.46
+(1e-3, 10, 500) 0.59 7.0e-3 488.73 697.33
+(1e-3, 10, 1000) 0.69 6.5e-3 967.56 1387.41
+(1e-1, 10, 2000) 0.73 6.0e-3 1921.29 2754.74
+ML-CW (0.01, 10, 1000) 0.64 1.4e-2 297.53 1679.72
+LOTS $2000$ 0.78 4.3e-3 242.23 600.14
+MLA-LP $1000$ 0.00 NA NA NA
+$3000$ 0.00 NA NA (>>1h)
+VII-C Multi-label classification (VOC2012 - InceptionV3)
+Table 11.
+shows the results in the case of multi-label classification when the target label vector is chosen randomly among those contained in the training dataset.
+The average number of labels targeted by the attacks is 2.1888Most of the images in the VOC2012 dataset contains one or very few labeled objects, hence the label vectors have few 1’s and choosing the target label from the training dataset results in few bit changes.
+It can be observed that both JMA and ML-C&W attack achieve ASR = 1.
+However, JMA has an advantage over ML-C&W both in terms of MSE and, most of all, in terms of computational complexity, being approximately 64 times faster than ML-C&W.
+LOTS achieves much worse performance, with ASR = 0.66 and a larger MSE.
 However, the average bASR is $0.979$, that confirms that the method has some effectiveness.
-We believe that the reason for the not so good results of LOTS relies on the construction of the target vector performed by the method.
-In fact, compared to multi-class classification where images belonging to the same class have similar content, e.g.
-images showing the same traffic sign, in the multi-label classification case, images sharing the same label vector may be very different from each other, and the differences between them grow with the amount of categories included in the label vector.
-For instance, an image of a crowded city street and an image with footballers on a soccer field are both instances of the people category, with the ’people’ label equal to 1.
+One possible interpretation for the not so good results of LOTS relies on the construction of the target vector performed by the method.
+In fact, compared to single-label classification, where images belonging to the same class have similar content (e.g.
+images showing the same traffic sign), in the multi-label classification case, images sharing the same label vector may be very different from each other.
+For instance, an image of a crowded city street and an image with footballers on a soccer field are both instances of the people category, with the ‘people’ label equal to 1.
 As LOTS chooses the target output vector by averaging the logits of the target images, the resulting target vector might not be plausible when the logits of the target images are significantly different.
-Finally, the performance of MLA-LP are also poor, with a much lower ASR w.r.t.
+Finally, the performance of MLA-LP and SemA-ML are also poor, with a much lower ASR w.r.t.
 JMA and ML-C&W.
-Moreover, the computational cost of the attack is very high.
-These results are not in contrast with the results reported in [11] for the same dataset, given that in [11] the method is validated for target attacks that change only one bit of the original label vector, while the average number of bits changed in our experiments is larger than 2.
-Tables VII through IX show the results in a challenging scenario where the target vector labels are not chosen from the training/test set but are obtained by randomly changing a certain number of bits of the label vector of the to-be-attacked image.
+Moreover, in the case of MLA-LP, the complexity of the attack is very high999 These results are not in contrast with the results reported in [10] for the same dataset, given that in [10] the method is validated for target attacks that change only one bit of the original label vector, while the average number of bits changed in our experiments is larger than 2..
+Tables VII through IX show the results in a challenging scenario where the target vector labels are not chosen from the training/test set but are obtained by randomly changing a certain number of bits of the groud truth label vector.
 Specifically, Tables VII, VIII and IX correspond to attacks inducing respectively 5, 10 and 20 bit errors (in the latter case all the bits should be changed).
-The results are reported for JMA, ML-C&W and MLA-LP.
-As we said LOTS can not be applied in this case, since this method requires the availability of a number of samples corresponding to the target label vector - that, in principle, in this case may not even correspond to a plausible image.
-Looking at the results in Table VII, we see that in the 5 bit case JMA can achieve ASR = 0.95 (bASR = 0.99), which is 5% higher than ML-CW with an average attack time with is approximately 6 times lower, while MLA-LP is totally ineffective (ASR = 0).
-For the case of 10 bit and 20 bit errors, we also report the results with larger $n_{it,max}$ for both JMA and ML-C&W.
-In fact, achieving a successful attack in the case where half of the bits and, even worse, all the bits are changed is extremely hard, and more iterations are necessary.
-From Table VIII we see that, in the 10 bit case, JMA can still achieve an ASR close to 0.90, at the price of a longer time necessary to run the attack, with an average number of iterations $\bar{n}_{it}$ that goes above 240.
-The C&W attack achieves a lower ASR with a bit larger PSNR when $n_{it,max}$ is raised to 5000, with a very large computational cost.
+The results are reported for JMA, ML-C&W, MLA-LP and SemA-ML.
+As stated previously, LOTS can not be applied in this case, since it requires the existence of a number of samples corresponding to the target label vector.
+Looking at the results in Table VII (5 flipping bits), JMA can achieve an ASR = 0.95 (bASR = 0.99), which is 5% higher than ML-CW, with an average attack time approximately 6 times faster, while MLA-LP and SemA-ML are totally ineffective (ASR = 0).
+As shown in Table VIII, in the 10 flipping bit case, JMA can still achieve an ASR close to 0.90, at the price of a longer time necessary to run the attack, with an average number of iterations $\bar{n}_{it}$ that goes above 240.
+C&W attack achieves a lower ASR when $n_{it,max}$ is raised to 5000, with a very large computational cost.
+The MSE reported for C&W is lower, however, this value is not computed on the same set of imagesand, arguably, a stronger distortion is necessary to attack the additional 16% images attacked by JMA.
 Fig.
 3 shows some examples of images successfully attacked by JMA in the 10-bit error case, when half of the labels are modified.
-As shown in Table IX, the ASR drops in the 20 bit case, namely, for an attack that tries to modify all the 20 components of the label vector.
+As shown in Table IX, the ASR drops in the 20 bit case (when the attacker tries to modify all the 20 bits of the label vector).
 Arguably, this represents a limit and extremely challenging scenario.
-We see that, by increasing $n_{it,max}$ to 2000, the JMA attack is successful 26% of the times, at the price of a high computational cost, while ML-C&W is never successful.
+It can be observed that, by increasing $n_{it,max}$ to 2000, JMA is successful 26% of the times, at the price of a high computational cost, while ML-C&W is never successful.
 Moreover, it is interesting to observe that JMA can achieve a bASR of 0.68, meaning that almost 70% of the labels are modified by the attack, while the bASR for ML-C&W is below 0.1.
-Not surprisingly, the PSNR is low in this case.
+Not surprisingly, the MSE is large in this case.
 Figure 3: Examples of JMA attacked images for which the attack can successfully flip 10 targeted bits of the label vector in the multi-label (VOC2012) classification task.
-For the car: the original image has label vector ( ’car’=1 , ‘person’ =0, ‘cat’ = 0, ‘cow’ = 0, ‘dog’ = 0, ‘sheep’ = 0, ‘bike’ = 0, ‘chair’ = 0, ‘potted plant’ = 0, ‘monitor’ = 0, ’aeroplane’ = 0, ‘bird’ = 0, ‘horse’ =0, ‘boat’ = 0, ‘dining table’ = 0, ’train’ = 0,…..), while the JMA attacked image is classified as ( ’car’=0 , ‘person’ =1, ‘cat’ = 1, ‘cow’ = 1, ‘dog’ = 1, ‘sheep’ = 1, ‘bike’ = 1, ‘chair’ = 1, ‘potted plant’ = 1, ‘monitor’ = 1, ’aeroplane’ = 0, ‘bird’ = 0, ‘horse’ =0, ‘boat’ = 0, ‘dining table’ = 0, ’train’ = 0,…..).
-For the airplane: the original image has label vector ( ’car’=0 , ‘person’ =0, ‘cat’ = 0, ‘cow’ = 0, ‘dog’ = 0, ‘sheep’ = 0, ‘bike’ = 0, ‘chair’ = 0, ‘potted plant’ = 0, ‘monitor’ = 0, ’aeroplane’ = 1, ‘bird’ = 0, ‘horse’ =0, ‘boat’ = 0, ‘dining table’ = 0, ’train’ = 0,…..), while the label vector of the JMA attacked image is ( ’car’=0 , ‘person’ =1, ‘cat’ = 1, ‘cow’ = 1, ‘dog’ = 0, ‘sheep’ = 0, ‘bike’ = 0, ‘chair’ = 0, ‘potted plant’ = 1, ‘monitor’ = 1, ’aeroplane’ = 0, ‘bird’ = 0, ‘horse’ =1, ‘boat’ = 1, ‘dining table’ = 1, ’train’ = 0,…..).
-The corresponding PSNR value is also reported.
-TABLE VI: Results of the attacks against multi-label classification on the VOC2012 dataset. On the average the attacks are asked to flip 2.1 labels.
-Parameters ASR bASR PSNR $\bar{n}_{it}$ Time(s)
-JMA (0.05,200) 1.00 1.0000 56.54 29.94 25.97
-(0.1,200) 1.00 1.0000 55.78 17.50 15.39
-(0.5,200) 1.00 1.0000 49.66 7.51 7.32
-(1,200) 1.00 1.0000 44.86 6.42 6.45
-ML-C&W (0.01, 10, 1000) 1.00 1.0000 52.03 180.93 469.90
-LOTS 2000 0.66 0.9787 50.91 437.20 268.16
-MLA-LP 100 0.58 0.9560 51.17 50.44 1361.29
+For the car: the original image has label vector (car=1, person=0, cat=0, cow=0, dog=0, sheep=0, bike=0, chair=0, potted-plant=0, monitor=0, aeroplane=0, bird=0, horse=0, boat=0, dining-table=0, train=0,...) while the JMA attacked image is classified as (car=0, person=1, cat=1, cow=1, dog=1, sheep=1, bike=1, chair=1, potted-plant=1, monitor=1, aeroplane=0, bird=0, horse=0, boat=0, dining-table=0, train=0, ...).
+For the airplane: the original image has label vector (car=0, person=0, cat=0, cow=0, dog=0, sheep=0, bike=0, chair=0, potted-plant=0, monitor=0, aeroplane=1, bird=0, horse=0, boat=0, dining-table=0, train=0,...), while the label vector of the JMA attacked image is (car=0, person=1, cat=1, cow=1, dog=0, sheep=0, bike=0, chair=0, potted-plant=1, monitor=1, aeroplane=0, bird=0, horse=1, boat=1, dining table=1, train=0,...).
+TABLE VI: Results against multi-label VOC2012 classification (the target label vector is taken from the training set).
+On average, the attacks are asked to flip 2.1 labels/bits.
+[^11] In this and the following tables, the parameter $n_{bs}$ is omitted for ML-CW, being always equal to 10.
+Parameters ASR bASR MSE $\bar{n}_{it}$ Time(s)
+JMA (0.05, 200) 1.00 1.0000 2.2e-6 29.94 25.97
+(0.1, 200) 1.00 1.0000 2.6e-6 17.50 15.39
+(0.5, 200) 1.00 1.0000 1.1e-5 7.51 7.32
+(1, 200) 1.00 1.0000 3.3e-5 6.42 6.45
+ML-C&W (0.01, 1000) 1.00 1.0000 6.3e-6 180.93 469.90
+LOTS 2000 0.66 0.9787 8.1e-6 437.20 268.16
+MLA-LP 100 0.58 0.9560 7.6e-6 50.44 1361.29
+SemA-ML 300 0.12 0.8742 8.02-5 30.30 6.55
 TABLE VII: Results for VOC2012 classification with random choice of the target label vector (5 bits out of 20 are flipped).
-Parameters ASR bASR PSNR $\bar{n}_{it}$ Time(s)
-JMA (0.05,200) 0.49 0.9032 49.18 129.85 139.31
-(0.1,200) 0.82 0.9687 47.48 110.21 106.98
-(0.2,200) 0.95 0.9942 45.26 87.96 79.05
-ML-C&W (0.01, 10, 1000) 0.90 0.9883 45.29 113.35 420.21
-MLA-LP 100 0 0.4264 NA NA NA
+Parameters ASR bASR MSE $\bar{n}_{it}$ Time(s)
+JMA (0.05, 200) 0.49 0.9032 1.2e-5 129.85 139.31
+(0.1, 200) 0.82 0.9687 1.8e-5 110.21 106.98
+(0.2, 200) 0.95 0.9942 3.0e-5 87.96 79.05
+ML-C&W (0.01, 1000) 0.90 0.9883 2.9e-5 113.35 420.21
+MLA-LP 100 0.00 0.4264 NA NA NA
+SemA-ML 300 0.00 0.7412 NA NA NA
 TABLE VIII: Results for VOC2012 classification with random choice of the target label vector (10 bits out of 20 are flipped).
-Parameters ASR bASR PSNR $\bar{n}_{it}$ Time(s)
-JMA (0.05,200) 0.01 0.5714 48.16 161.00 157.41
-(0.1,200) 0.06 0.6700 45.62 163.84 156.04
-(0.2,200) 0.23 0.8112 42.77 163.40 162.63
-(0.5,200) 0.31 0.8842 35.60 161.36 158.88
-(0.1, 500) 0.56 0.8892 43.61 312.42 397.42
-(0.5, 500) 0.88 0.9805 35.66 245.24 278.02
-ML-C&W (0.01, 10, 1000) 0.295 0.8697 45.09 270.17 498.63
-(0.01, 10, 2000) 0.4186 0.9156 44.46 451.78 939.53
-(0.1, 10, 2000) 0.475 0.9147 44.73 433.75 934.58
-(0.5, 10, 2000) 0.475 0.9147 44.78 430.96 908.92
-(0.1, 10, 3000) 0.575 0.9331 44.81 581.53 1461.05
-(0.1, 10, 5000) 0.72 0.9557 44.74 775.97 2183.79
-MLA-LP 100 0 0.5167 NA NA NA
+Parameters ASR bASR MSE $\bar{n}_{it}$ Time(s)
+JMA (0.05,200) 0.01 0.5714 1.5e-5 161.00 157.41
+(0.1, 500) 0.56 0.8892 4.3e-5 312.42 397.42
+(0.5, 500) 0.88 0.9805 2.7e-4 245.24 278.02
+ML-C&W (0.01, 1000) 0.295 0.8697 3.2e-5 270.17 498.63
+(0.01, 2000) 0.4186 0.9156 3.6e-5 451.78 939.53
+(0.1, 2000) 0.475 0.9147 3.1e-5 433.75 934.58
+(0.5, 2000) 0.475 0.9147 3.3e-5 430.96 908.92
+(0.1, 3000) 0.575 0.9331 3.3e-5 581.53 1461.05
+(0.1, 5000) 0.72 0.9557 3.4e-5 775.97 2183.79
+MLA-LP 100 0.00 0.5167 NA NA NA
+SemA-ML 300 0.00 0.4994 NA NA NA
 TABLE IX: Results for VOC2012 classification with random choice of the target label vector. Case of 20 bit errors (all labels are changed by the attack).
-Parameters ASR bASR PSNR $\bar{n}_{it}$ Time(s)
+Parameters ASR bASR MSE $\bar{n}_{it}$ Time(s)
 JMA (0.05, 200) 0 0.1109 NA NA NA
@@
 (0.5, 500) 0 0.3288 NA NA NA
-(0.5,1000) 0.03 0.5068 30.86 938.16 980.82
-(0.5,2000) 0.26 0.6795 29.43 1484.34 1677.49
+(0.5, 1000) 0.03 0.5068 8.2e-4 938.16 980.82
+(0.5, 2000) 0.26 0.6795 1.1e-3 1484.34 1677.49
 ML-C&W (0.01, 10, 1000) 0 0.0659 NA NA NA
 (0.5, 10, 2000) 0 0.0882 NA NA NA
-7 Discussion
-Our experiments showed that JMA requires a small number of iterations to find an adversarial example, much smaller than the other algorithms, and gets an ASR that is generally higher (in some cases, much higher) for a similar distortion.
-It is worth stressing that the number of updates $\bar{n}_{it}$ reported in the tables takes into account also the steps of the binary search carried out at the end of the iterations (see Section 5.4).
-Specifically, $it=(v-1)+n_{bs}$, where $v$ is the number of ’for’ loops, that is, the number of image updates (see Algorithm 1), and $n_{bs}$ is set equal to $6$.
-Then, the $\bar{n}_{it}$ values reported in the various tables reveal that JMA can indeed find an adversarial image in one shot in many cases, see for instance Table IV for CIFAR-10 ECOC classification and Table VI for the multi-label classification where $\bar{n}_{it}$ is lower than 7 for the attack settings with larger $\epsilon$ [^10] Note that, even when a one-shot attack is possible, a perturbation applied to the image with a weak strength $\epsilon$ may not result in an adversarial image, thus requiring that more iterations are run..
+TABLE X: Results for VOC2012 using different architectures (target label vector from the training set): (a) ResNet50, (b) ViT-B/16 and (c) CLIP+MLP.
+Parameters ASR bASR MSE SSIM $\bar{n}_{it}$ Time(s)
+JMA (0.05, 200) 1.00 1.0000 3.3e-6 0.999 32.50 2.88
+(0.1, 200) 1.00 1.0000 3.8e-6 0.999 21.96 1.60
+(0.5, 200) 1.00 1.0000 1.4e-5 0.993 12.84 0.79
+(1, 200) 1.00 1.0000 4.0e-5 0.982 11.84 0.80
+ML-C&W (0.01,1000) 0.58 0.9583 3.1e-5 0.989 761.00 307.27
+LOTS 2000 0.83 0.9610 2.0e-2 0.386 34.84 138.20
+MLA-LP 100 0.22 0.9500 3.5e-4 0.964 10.98 648.39
+SemA-ML 300 0.31 0.9167 9.3e-5 0.969 5.30 5.45
+(a) ResNet50
+Parameters ASR bASR MSE SSIM $\bar{n}_{it}$ Time(s)
+JMA (0.05, 200) 0.99 0.9993 4.4e-5 0.990 26.01 3.89
+(0.1, 200) 1.00 1.0000 4.4e-5 0.990 17.65 2.22
+(0.5, 200) 1.00 1.0000 3.9e-5 0.991 9.49 0.85
+(1, 200) 1.00 1.0000 4.6e-5 0.989 8.48 0.69
+ML-C&W (0.01, 1000) 0.81 0.9825 3.3e-5 0.992 635.34 996.74
+LOTS 2000 0.84 0.9750 1.5e-2 0.603 38.00 217.00
+MLA-LP 100 0.52 0.9705 6.0e-4 0.910 5.79 787.87
+SemA-ML 300 0.16 0.8835 8.6e-5 0.974 17.91 11.52
+(b) ViT-B-16
+Parameters ASR bASR MSE SSIM $\bar{n}_{it}$ Time(s)
+JMA (0.05, 200) 1.00 1.0000 1.6e-6 0.999 28.65 4.46
+(0.1, 200) 1.00 1.0000 2.3e-6 0.999 19.72 2.81
+(0.5, 200) 1.00 1.0000 1.2e-5 0.994 11.88 1.30
+(1, 200) 1.00 1.0000 3.5e-5 0.984 11.07 1.13
+ML-C&W (0.01, 1000) 0.14 0.9095 4.8E-05 0.983 684.75 1613.82
+LOTS 2000 0.97 0.9818 1.3E-02 0.496 19.59 106.17
+MLA-LP 100 0.11 0.9400 2.1E-04 0.977 5.57 710.85
+SemA-ML 300 0.09 0.8788 9.9E-05 0.973 1.00 21.84
+(c) CLIP+MLP
+TABLE XI: Results for VOC2012, for a 10-bit flipping attack (10 bits out of 20 are flipped). (a) ResNet50, (b) ViT-B/16 and (c) CLIP+MLP.
+Parameters ASR bASR MSE SSIM $\bar{n}_{it}$ Time(s)
+JMA (0.05, 200) 0.00 0.5483 NA NA NA NA
+(0.1, 500) 0.94 0.9780 4.8e-5 0.984 291.08 25.76
+(0.5, 500) 1.00 1.0000 2.4e-4 0.916 166.60 14.66
+ML-C&W (0.01, 1000) 0.37 0.8805 5.0e-4 0.884 631.78 275.72
+(0.5, 2000) 0.59 0.8808 1.1e-1 0.155 1108.74 1154.25
+(0.1, 3000) 0.42 0.8710 1.6e-2 0.464 581.57 1747.41
+(0.1, 5000) 0.45 0.8690 1.6e-2 0.470 756.03 3340.60
+MLA-LP 100 0.00 0.5625 NA NA NA NA
+SemA-ML 300 0.00 0.4980 NA NA NA NA
+(a) ResNet50
+Parameters ASR bASR MSE SSIM $\bar{n}_{it}$ Time(s)
+JMA (0.05, 200) 0.65 0.9073 1.1e-4 0.974 154.64 25.00
+(0.1, 500) 1.00 1.0000 1.6e-4 0.967 126.64 19.87
+(0.5, 500) 1.00 1.0000 2.6e-4 0.948 63.89 9.79
+ML-C&W (0.01, 1000) 0.51 0.8680 5.5e-4 0.903 582.63 996.45
+(0.5, 2000) 0.74 0.8888 5.6e-2 0.272 358.84 3230.31
+(0.1, 3000) 0.62 0.8445 1.2e-2 0.584 912.28 4074.83
+(0.1, 5000) 0.66 0.8445 1.9e-2 0.565 1497.76 6674.43
+MLA-LP 100 0.00 0.5635 NA NA NA NA
+SemA-ML 300 0.00 0.4995 NA NA NA NA
+(b) ViT-B/16
+Parameters ASR bASR MSE SSIM $\bar{n}_{it}$ Time(s)
+JMA (0.05, 200) 1.00 0.9975 3.8e-6 0.998 59.16 8.70
+(0.1, 500) 1.00 1.0000 5.8e-6 0.997 41.45 5.86
+(0.5, 500) 1.00 1.0000 3.3e-5 0.986 22.73 2.79
+ML-C&W (0.01, 1000) 0.51 0.8680 5.5e-4 0.903 582.63 2052.44
+(0.5, 2000) 0.74 0.8888 5.6e-2 0.272 358.84 3600.40
+(0.1, 3000) 0.62 0.8445 1.2e-2 0.584 912.28 4149.43
+(0.1, 5000) 0.66 0.8445 1.9e-2 0.565 1497.76 4867.13
+MLA-LP 100 0.00 0.5755 NA NA NA NA
+SemA-ML 300 0.00 0.4970 NA NA NA NA
+(c) CLIP+MLP
+VII-D Multi-label classification (VOC2012 - ResNet50, ViT-B/16, CLIP+MLP)
+To validate the effectiveness of JMA against different types of architectures, we also trained a residual network (ResNet50), a Transformer-based architecture (ViT-B/16) and a network based on pre-trained vision-language models (CLIP), obtained by adding a lightweight classification head - namely a Multi-Layer Perceptron - on top of CLIP features (CLIP+MLP).
+The models were trained using the multi-label cross-entropy loss with Adam optimizer with lr = $10^{-3}$, batch size of 32 samples, and early stopping based on the mAP computed on the validation set.
+The ResNet50, ViT-B/16, and CLIP+MLP models obtained a mAP score on the test set of 0.94, 0.87, and 0.90, respectively.
+Table X and XI report the results respectively in the case of attack target vector chosen randomly from the training set and in the case where 10 bit at random are modified in the label vector.
+For a more comprehensive evaluation on the quality of attacked images and the amount of distortion introduced by the attack, in the tables we also report the SSIM in addition to MSE.
+We see that the state-of-the-art methods, especially ML-C&W and MLA-LP, achieve worse performance than in the InceptionV3 case, while LOTS is more effective.
+With regard to JMA, performance are very good and ASR = 1 can be obtained in all the cases and in both attack settings.
+Furthermore, JMA is the attack introducing the lowest distortion, with an MSE in the range $10^{-6},10^{-4}]$.
+The SSIM is also very good, being always above 0.99 in the first attack setting, and remaining good also in the challenging 10-bit flips attack.
+Compared to the InceptionV3 case, JMA takes more iterations to run ($\bar{n}_{it}$ is larger), however the algorithm is extremely fast and the attacked image can be found in a few seconds in most cases.
+TABLE XII: Results of the 10-bit flipping attack on MS-COCO dataset.
+Parameters ASR bASR MSE SSIM $\bar{n}_{it}$ Time(s)
+JMA (0.05,200) 1.00 1.0000 7.7e-6 0.997 53.80 76.86
+(0.1, 500) 1.00 1.0000 9.5e-6 0.997 38.59 51.66
+(0.5, 500) 1.00 1.0000 4.4e-5 0.984 30.25 40.24
+ML-C&W (0.01, 1000) 0.00 0.7649 NA NA NA NA
+(0.5, 2000) 0.00 0.7984 NA NA NA NA
+(0.1, 3000) 0.00 0.7578 NA NA NA NA
+MLA-LP 100 0.00 0.8892 NA NA NA NA
+SemA-ML 300 0.00 0.8518 NA NA NA NA
+(a) Swin-B
+Parameters ASR bASR MSE SSIM $\bar{n}_{it}$ Time(s)
+JMA (0.05, 200) 0.46 0.9177 8.9e-6 0.996 149.17 98.73
+(0.1, 500) 1.00 1.000 1.9e-5 0.993 163.08 109.78
+(0.5, 500) 1.00 1.000 1.4e-4 0.957 101.66 66.38
+ML-C&W (0.01, 1000) 0.21 0.9373 5.3e-4 0.905 320.93 1377.91
+(0.5, 2000) 0.11 0.9476 1.1e-1 0.222 853.68 3154.67
+(0.1, 3000) 0.19 0.9331 2.1e-2 0.481 793.53 4507.12
+MLA-LP 100 0.00 0.8816 NA NA NA NA
+SemA-ML 300 0.00 0.8596 NA NA NA NA
+(b) CLIP+MLP
+TABLE XIII: Results of the 10-bit flipping attack on NUS-WIDE dataset (the network is CLIP+MLP).
+Parameters ASR bASR MSE SSIM $\bar{n}_{it}$ Time(s)
+JMA (0.05,200) 0.99 0.9975 5.1e-6 0.998 81.65 48.80
+(0.1, 500) 1.00 1.00 7.8e-6 0.996 59.96 35.23
+(0.5, 500) 1.00 1.00 5.3e-5 0.976 36.53 20.35
+ML-C&W (0.01, 1000) 0.00 0.7559 NA NA NA NA
+(0.5, 2000) 0.00 0.7612 NA NA NA NA
+(0.1, 3000) 0.00 0.7070 NA NA NA NA
+MLA-LP 100 0.00 0.8843 NA NA NA NA
+SemA-ML 300 0.00 0.8630 NA NA NA NA
+TABLE XIV: Comparison between JMA and SemA-ML [11] on the NUS-WIDE dataset in the case of semantic consistent multi-label attacks.
+Parameters ASR bASR MSE SSIM $\bar{n}_{it}$ Time(s)
+JMA (0.05, 200) 1.00 1.0000 1.2e-6 0.999 26.98 14.16
+(0.1, 500) 1.00 1.0000 1.8e-6 0.999 19.57 9.39
+(0.5, 500) 1.00 1.0000 1.2e-5 0.994 12.33 4.76
+SemA-ML 300 0.20 0.9568 7.2e-5 0.978 36.43 5.10
+(a) Target label vector from training dataset. The average number of bits that the attack is required to flip is 4.37 bit.
+Parameters ASR bASR MSE SSIM $\bar{n}_{it}$ Time(s)
+JMA (0.05, 200) 1.00 1.0000 6.6e-7 1.000 19.03 8.84
+(0.1, 500) 1.00 1.0000 1.0e-6 0.999 14.17 5.77
+(0.5, 500) 1.00 1.0000 6.6e-6 0.996 9.58 2.88
+SemA-ML 300 0.48 0.9709 5.8e-5 0.975 13.65 4.20
+(b) Attack removing all labels/bits. The average number of bits that the attack is required to flip is 2.90 bit.
+VII-E Multi-label classification (MS-COCO, NUS-WIDE)
+To validate the effectiveness of JMA against more challenging datasets, we considered the MS-COCO [41] and NUS-WIDE [42] datasets, containing 80 and 81 labels respectively, resulting in a huge number of possible label vectors.
+For the MS-COCO dataset, we trained a Swin-B architecture (which obtained a better mAP than the ViT-B/16) and CLIP+MLP, with input size of $224\times 224$ pixels.
+The models were trained on the original split of the dataset using the Asymmetric Loss (ASL) proposed in [43], with Adam optimizer, lr = $10^{-3}$ and batch size of 32 samples.
+Early stopping on the validation set was employed based on mAP.
+The Swin-B and CLIP+MLP models scored an mAP of about 0.82 and 0.81, which is aligned with benchmarks for similar input shapes.
+For the NUS-WIDE case, we split the dataset into 129,431 training, 32,358 validation, and 107,759 test images.
+A CLIP+MLP architecture was trained, using the same learning setting as for MS-COCO, obtaining a mAP score of 0.59, which is aligned with banchmarks.
+Table XII reports the results on MS-COCO in the 10-bit flipping attack scenario.
+We see that the architecture has a noticeable impact on the performance of the attacks.
+For the state-of-the-art methods, attacking the Swin-B network is more difficult.
+Specifically, the bASR values achieved by ML-C&W range from 0.75 to 0.80 and from 0.93 to 0.95 across the various parameter settings respectively in the Swin-B and CLIP+MLP case, thus being similar to the VOC2012 case.
+However, the ASR is much lower and only very few images can be fully attacked in the CLIP+MLP case, while in the case of Swin-B the ASR is 0.
+On the contrary, MLA-LP and SemA-ML work better on MS-COCO w.r.t.
+the VOC2012 case and the bASR is higher.
+Regarding JMA, in both cases the attack can reach ASR = 1 with low attack distortion, thus surpassing the state-of-the-art.
+In the CLIP+MLP case, the attack takes more time to run and the distortion introduced by the attack is larger, yet the SSIM remains above 0.99.
+For the NUS-WIDE dataset, the results of the 10-bit flipping attack are reported in Table XIII.
+We see that the ASR is always 0 for the state-of-the-art methods, while attacking with JMA resulted in ASR = 1.
+Results in terms of distortion and computational efficiency are similar to the best results on MS-COCO, achieved using Swin-B.
+It is worth observing that SemA-ML is a multi-label attack which enforces that semantic consistency among the labels is retained by the attack.
+Differently from the case of VOC2012 and MS-COCO, in the NUS-WIDE dataset, labels correspond to both categories and subcategories (e.g.
+, ‘vehicle’ and ‘car’ are distinct labels).
+Hence, semantic consistence may not be verified when bits are flipped at random (e.g., the label ‘car’ may be turned on while the label ‘vehicle’ remains off, leading to a semantically inconsistent vector).
+Hence, to compare JMA with SemA-ML in a setting that satisfies this method’s working assumptions, we also ran some tests considering the following simple attack settings, where semantic consistency is satisfied: i) target label vector chosen at random from the training/test set; ii) target vector with all labels/bits set to zero.
+The results in these cases are reported in Table XIV.
+SemA-ML can now attack most of the labels and reach the target in some cases (and the ASR is no longer 0).
+This is especially the case when the attack is asked to set all the bits to zero (in this case, the attack involves flipping a lower number of bits - 2.90 on the average).
+In this case, SemA-ML can achieve bASR = 0.971 and ASR = 0.48.
+In any case, the performance of JMA remains largely superior, achieving ASR = 1 in all the settings with a lower MSE and an SSIM above 0.99 (still with a lower complexity).
+VII-F JMA as a one-step attack
+Experiments show that JMA requires a small number of iterations to find an adversarial example, much smaller than the other algorithms, and gets an ASR that is generally higher (in some cases, much higher) for a similar distortion.
+It is worth stressing that the number of updates $\bar{n}_{it}$ reported in the tables takes into account also the steps of the binary search carried out at the end of the iterations (see Section VI-D).
+Specifically, $it=(v-1)+n_{bs}$, where $v$ is the number of ‘for’ loops, that is, the number of image updates (see Algorithm 1), and $n_{bs}$ is set equal to $6$.
+Therefore, JMA can indeed find an adversarial image in one shot in a considerable number of cases, see for instance Table IV for CIFAR-10 ECOC classification and Table 11 for the multi-label classification, where $\bar{n}_{it}$ is lower than 7 for the attack settings with larger $\epsilon$ [^12] Note that, even when a one-shot attack is possible, a perturbation applied to the image with a weak strength $\epsilon$ may not result in an adversarial image, thus requiring that more iterations are run..
 This confirms a-posteriori the validity of one of the most basic assumptions underlying our approach, namely the local linear behavior, that, when it holds, allows to perform the attack in one-shot.
-When the linear approximation does not hold in the close vicinity of the initial point, and possibly of other points obtained after the updating, the perturbation is added with a small $\epsilon$ in order to change the initial point, and the Jacobian is recomputed (see Algorithm 1).
-By inspecting the tables (see for instance Table II), we can see that, when this happens, it is often preferable to use a not-too-small $\epsilon$ so that the point is moved farther from the initial point, where the behavior of the local function can be closer to linear, instead of remaining in the vicinity of the initial point where the local approximation may still be not satisfied.
-Doing so, a high ASR can be achieved with a reduced number of iterations and a similar PSNR.
-8 Conclusions
+When the linear approximation does not hold in the close vicinity of the initial point, and possibly of other points during the updates, the perturbation is added with a small $\epsilon$ in order to change the initial point, and the Jacobian is recomputed (see Algorithm 1).
+By inspecting the tables (see for instance Table I), we can see that, when this happens, it is often preferable to use a not-too-small $\epsilon$ so that the point is moved farther from the initial point, where the behavior of the local function can be closer to linear, instead of remaining in the vicinity of the initial point where the local approximation may still be not satisfied.
+Doing so, a high ASR can be achieved with a reduced number of iterations and a similar distortion.
+VIII Concluding remarks
 We have proposed a general, nearly optimal, targeted attack, that can solve the original formulation of the adversarial attack by Szegedy et al.
 under a first order approximation of the network function.
-The method resorts to the minimization of a Mahalanobis distance term, which depends on the Jacobian matrix taking into account the effort necessary to move the output representation of the image in a given direction.
+The method resorts to the minimization of a Mahalanobis distance term, which depends on the Jacobian matrix taking into account the effort necessary to move the feature representation of the image in a given direction.
 By exploiting the Wolfe duality, the minimization problem is reduced to a non-negative least square (NNLS) problem, that can be solved numerically.
-The experiments show that the JMA attack is effective against a wide variety of DNNs adopting different output encoding schemes, including networks using error correction output encoding (ECOC) and multi-label classification, outperforming existing attacks in terms of larger attack success rate, lower distortion and lower complexity.
-In particular, in the multi-label case, our experiments show that JMA is capable to simultaneously change up to 10 (random) labels of the output label vector for VOC2012 classification with a very high ASR (and, in some cases, it can also change all the 20 output labels), a capability that is completely out of reach of the attacks proposed so far.
-JMA remains effective also in the case of one-hot encoding, where it has much reduced computational complexity with respect to the C&W attack.
-Future work will focus on exploiting the JMA attack in a black-box attack scenario, to develop powerful targeted attacks with certain transferability properties.
-Exploiting the proposed attack for adversarial training, to increase robustness of general classification models against targeted attacks, is also another interesting research direction.
-Acknowledgment
-This work has been partially supported by the Italian Ministry of University and Research under contract PRIN 2017 2017Z595XS-001 (PREMIER project).
+The experiments show that the JMA attack is effective against a wide variety of DNNs adopting different output encoding schemes, including networks using error correction output encoding (ECOC) and in particular multi-label classification, outperforming existing attacks in terms of higher ASR, lower distortion and lower complexity, with attack capabilities far exceeding those of existing attacks.
+JMA remains effective also in the case of one-hot encoding, with much reduced computational complexity with respect for instance to the C&W attack.
+Finally, we observe that JMA can inspire new defences.
+In particular, given its efficiency, JMA can be used to reduce the load of adversarial training and develop effective adversarial training defences [2, 44, 45].
+Furthermore, while with the exception of scattered works ([44]) adversarial training is primarly limited to applications of single-label classification, JMA can be used to implement effective adversarial training also for other types of classifiers (e.g.
+multi-label classifiers).
+In addition, thanks to its efficiency, JMA could also be used to perform adversarial training of provably robust multi-label classifiers exploiting randomized smoothing [46].
+In fact, similarly to what has been done for standard single-label randomized smoothing classifiers ([47]), by training with JMA attacked samples it could be possible to boost the provable robustness of smoothed multi-label classifiers.
+Future work could also focus on the extension of JMA to a black-box attack scenario, to develop powerful targeted attacks with certain transferability properties.